### v2

# DYNAMICAL HORIZONS AND BLACK HOLE SOFT HAIR

Abstract
In the present work, quasilocal Brown-York charges are derived that coincide in the large sphere limit with the conserved supertranslation hair and superrotation charges introduced by Hawking, Perry and Strominger in [45, 46]. Given these charges, a general scenario is outlined in which a non-rotating black hole completely evaporates after its collapse due to particle creation effects, whereby a genuine one-way traversable event horizon is never formed, but merely a two-way traversable dynamical (resp. future trapping) horizon. The formation of such a dynamical horizon has the consequence, as is demonstrated, that quasilocal energy transported by the considered charges, and thus information, can continuously escape through the black hole horizon to infinity; a mechanism which, as is argued, could possibly prevent information loss once the black hole formation and evaporation process comes to an end.
Key words: Black hole soft hair, dynamical horizons, Hawking effect;
Introduction
It is generally expected that black holes can evaporate due to contiuous emission of thermal radiation. Responsible for this development are particle creation effects first discovered by Hawking almost half a century ago [43], whose existence has led to one of the most persistent problems of modern theoretical physics: the black hole information loss paradox.
This paradox occurs naturally when considering quantized matter collapsing into a black hole that evaporates in finite time. More precisely, when considering, a la Hawking, a quantized massless scalar field in an approximately classical geometric background field in the context of semiclassical Euclidean quantum gravity, this paradox arises naturally from the fact that quantum correlations overcoming (by tunneling through) the gravitational potential barrier of the black hole cause a continuous emission of radiative modes to infinity. This has the consequence that the asymptotic vacua at past and future null infinity do not coincide, thereby implying that the past ground state of the field must be a thermal state with respect to the future ground state and vice versa. The latter is related to the fact that certain quantum correlations are shielded by the black hole singularity and therefore cannot be restored at future null infinity at the end of the evaporation process. This, however, allows for the conclusion that the evolution from low-entropy collapsing matter to high-entropy radiation must result in a loss of quantum information.
Notwithstanding the fact that the validity of these results of Hawking’s work have not been questioned in any of the relevant literature on the subject, this outcome of the process is generally deemed physically implausible. After all, it would imply that black hole formation and evaporation processes would inevitably lead to a breakdown of predictability at the quantum level [44, 47].
In response to this obstacle, numerous approaches have been pursued in the literature to solve the information loss problem and show that quantum information is preserved in black hole evaporation processes; see here e.g. [50, 61, 62, 73, 74, 77] and references therein for an overview of the subject.
Among the approaches mentioned, there is a fairly recent proposal focusing on shock-wave-induced black hole soft hair, i.e. supertranslation and/or superrotation charges conserved by asymptotic Bondi-Metzner-Sachs (BMS) symmetries in asymptotically flat spacetimes. First considered in [32, 45, 46], the existence of such conserved charges remarkably does not contradict the no-hair theorem; for the simple reason that the supertranslated black hole geometries considered are diffeomorphic to untranslated ones and thus exhibit geometric properties that do not conflict with the classic laws of black hole physics. In addition, the BMS invariance of said charges suggests that there should be an infinite family of degenerate vacua related by infinite-dimensional asymptotic symmetries and thus an infinite number of conservation laws governing gravitational scattering processes at null infinity. In light of these facts, the authors of [45, 46] expressed the idea that black hole soft hairs could encode the missing part of the information that is lost during black hole formation and evaporation process; thus possibly contributing to the solution of the information paradox.
Yet, reservations have been expressed about this proposal [20, 21, 63] for two reasons: $i)$ soft hair should not be able to store information and $ii)$ the arising quantum correlations should be too small to purify Hawking radiation.
What is generally expected to be true, however, is that soft hair leads to measurable effects such as the gravitational memory and spin memory effects; as well as non-trivial quantum effects, where the latter are likely to be of relevance to black hole physics [41, 67, 72]. Still, the majority of approaches set out in the literature still treat soft hair exclusively classically without taking these very quantum effects into account.
Against this background, this paper addresses two main objectives:
The first one is to generalize the results of [53, 54] and show that, within the quasilocal Brown-York framework [16, 22, 23, 24], integral expressions can be defined which asymptotically reduce - as shown in section one and Appendix A of this work - to the conserved supertranslation hair and superrotation charges derived in [45, 46]. For this very purpose, a geometric framework is introduced that comprises a spatially and temporally bounded spacetime with inner and outer boundaries, where the inner boundary is formed by a dynamical black hole horizon in the sense of Ashtekar et al. [6, 7]. Taking this framework as a starting point, a null geometric approach is pursued, which serves as the basis for calculating the flux of mass and/or radiant energy through said dynamical horizon in a non-stationary black hole spacetime by performing a variation of the total gravitational Hamiltonian of the theory (bulk plus boundary term). The latter then leads to the aforementioned BMS charges and the associated black hole soft hairs when the outer boundary of spacetime is shifted to infinity in the large sphere limit.
The second objective, on the other hand, is to incorporate particle creation effects into the formalism used. To this end, the classic case of a quantized massless scalar field in a collapsing black hole background is considered within the semiclassical approximation of Euclidean quantum gravity. For said case, it is shown that boundary charges can be defined that asymptotically reduce to BMS charges in Bondi coordinates near null infinity; using in this context a semiclassical (intergral law) version of the null Raychaudhuri equation derived in section two of this work. Based on this result, the obtained expressions are shown to allow for the consideration of quantum backreaction phenomena caused by particle creation and the underlying Hawking effect; thus describing, as is shown, ’quantum analogs’ of conserved boundary charges at the black hole horizon that naturally soften and take the form of black hole supertranslation hairs as they approach past and/or future null infinity in the large sphere limit. These quantum charges are constructed using a semiclassical extension of the Bondi-Sachs formalism briefly discussed in Appendix A of this paper.
Given these insights, a general scenario is outlined in the rear part of the second section of the paper in which a black hole first forms and then completely evaporates, whereby the event horizon - in the spirit of earlier works by Hayward and Ashtekar [1, 49] - is replaced by a dynamical (resp. future trapping) horizon. Using a model of a Vaidya-type sandwich spacetime presented in Appendix B as an inspiration, it is then made clear that, according to the scenario described, it should be possible for a flux of quasilocal mass and/or radiant energy to escape through the dynamical horizon of an evaporating black hole to future null infinity; where the quasilocal charges responsible for this energy transfer are identified as those derived in the paper, generalizing the supertranslation charges of Hawking, Perry and Strominger. Eventually, by imposing conditions on the mode functions of the scalar field, again derived with respect to the toy model used, new arguments are provided for the conclusion that particle creation effects in non-stationary black hole spacetimes with locally defined dynamical horizons cannot cause information loss at the quantum level.
Consequences of these discoveries are discussed towards the end of the paper.
1 Quasilocal Charges in Bounded and Unbounded Gravitational Fields
In this section, using a null geometric approach to the quasilocal Brown-York framework [16, 22, 23, 24, 54], an expression is derived for the mass and/or radiative energy flux through a two-surface enclosing a nonstationary black hole spacetime. To this avail, a system of integral expressions is defined that reduces in an asymptically flat black hole background in the Bondi gauge close to null infinity to the conserved supertranslation hair and superrotation charges derived in [32, 46]. Moreover, antipodal matching conditions, first introduced in [46], are discussed and implications of these conditions for the black hole information loss paradox are pointed out. Eventually, following the reasoning of [54], generalized results on mass and radiative transfer in bounded and unbounded gravitational fields are derived, which lead to quasilocal corrections to the Bondi mass-loss formula and thus make clear that even more general boundary charges can be constructed by the approach to be introduced.
1.1 Geometric Framework
[…]
To further explain the notation, it may be noted that the the induced three-metric at $\Sigma_{t}$ is $h_{ab}=g_{ab}+n_{a}n_{b}$. The induced three-metric at $\mathcal{B}$, one the other hand, takes the form $\gamma_{ab}=g_{ab}-u_{a}u_{b}$ , with $u^{a}$ denoting a spatial unit vector field which is perpendicular to $\mathcal{B}$ and orthogonal to an associated temporal unit vector field $v^{a}$ tangent to $\mathcal{B}$. Moreover, in considering a further spacelike vector field $s^{a}$ that is orthogonal to the timelike generator $n^{a}$ of the spacelike foliation $\{\Sigma_{t}\}$, but generally non-orthogonal to $v^{a}$ (in contrast to $u^{a}$, which is generally non-orthogonal to $n^{a}$), one finds that the induced two-metric $q_{ab}$ at $\Omega_{t}$ takes the form $q_{ab}=g_{ab}-u_{a}u_{b}+v_{a}v_{b}=g_{ab}+n_{a}n_{b}-s_{a}s_{b}$. Using the latter relation in combination with the decompositions $n^{a}=\frac{1}{\sqrt{2}}(k^{a}+l^{a})$ and $s^{a}=\frac{1}{\sqrt{2}}(k^{a}-l^{a})$ of $n^{a}$ and $s^{a}$, where $l^{a}$ and $k^{a}$ are null normals reducing locally to those associated with a given leaf $\mathcal{S}_{t}$ of the foliation $\{\mathcal{S}_{t}\}$ of the dynamical horizon $\mathcal{T}$, one then finds that the induced metric at said horizon takes the previously claimed form $q_{ab}=g_{ab}+l_{a}k_{b}+k_{a}l_{b}$. Eventually, with respect to this induced metric, the boundary shift vector reads $\mathcal{N}^{a}=q_{\;c}^{a}N^{c}$.
Figure 1: A schematic three-dimensional representation of the spacetime manifold $M$ along with its boundaries.
1.2 Quasilocal Brown-York Charges, Variation of the Gravitational Hamiltonian
The starting point for all considerations to be made in this subsection is the total gravitational Hamiltonian of the theory. Given a spacelike hypersurface $\mathcal{C}$, this Hamiltonian is found to consist of a bulk and a boundary part, meaning that $H=H_{Bulk}+H_{Boundary}$, where $H_{Bulk}=\underset{\mathcal{C}}{\int}\mathcal{H}(N,\vec{N})d^{3}x$ and $H_{Boundary}=-\underset{\partial\mathcal{C}}{\oint}\mathfrak{h}(\mathcal{N},\vec{\mathcal{N}})d^{2}x$ are the corresponding bulk and boundary parts; thus giving rise to the integral expression
$H(N,\vec{N};\mathcal{N},\vec{\mathcal{N}})=\underset{\mathcal{C}}{\int}\mathscr{H}(N,\vec{N})d^{3}x-\underset{\partial\mathcal{C}}{\oint}\mathfrak{H}(\mathcal{N},\vec{\mathcal{N}})d^{2}x.$ (1)
Considering here the fact that the boundary of spacetime consists of two individual parts, it is found that the Hamiltonian (1)(1)( 1 ) splits in two different integral expressions from now on two be referred as interior and exterior parts of the Hamiltonian, i.e.
$H_{0}^{Bulk}(N,\vec{N})=$ $\underset{\Sigma_{t}}{\int}\mathscr{H}(N,\vec{N})d^{3}x,\;H_{0}^{Boundary}(\mathcal{N},\vec{\mathcal{N}})=\underset{\Omega_{t}}{\int}\mathfrak{H}(\mathcal{N},\vec{\mathcal{N}})d^{2}x$ (2)
$H_{h}^{Bulk}(N,\vec{N})=$ $\underset{\mathcal{H}}{\int}\mathscr{H}(N,\vec{N})d^{3}x,\;H_{h}^{Boundary}(\mathcal{N},\vec{\mathcal{N}})=\underset{\mathcal{S}_{t}}{\int}\mathfrak{H}(\mathcal{N},\vec{\mathcal{N}})d^{2}x,$
where $\mathscr{H}(N,\vec{N})$ represents the ADM Hamiltonian density $\mathscr{H}=\frac{\sqrt{h}}{8\pi}(G_{ab}-8\pi T_{ab})t^{a}n^{b}=\frac{\sqrt{h}}{8\pi}(N\mathcal{H}+\mathcal{H}_{a}N^{a})$, leading one to conclude that the bulk part $H_{Bulk}$ of the Hamiltonian (1)(1)( 1 ) vanishes on-shell. The local Hamiltonian density $\mathfrak{H}(\mathcal{N},\vec{\mathcal{N}})$ at $\Omega_{t}$, on the other hand, can be written in the form $\mathfrak{H}=\frac{\sqrt{q}}{8\pi}\rho_{ab}t^{a}v^{b}=\frac{\sqrt{q}}{8\pi}(\mathcal{N}\mathfrak{h}+\mathfrak{h}_{a}\mathcal{N}^{a})$, by virtue of the fact that the definitions $\mathfrak{h}=\frac{\sqrt{q}}{8\pi}\mathfrak{K}$ and $\mathfrak{h}_{a}=\frac{\sqrt{q}}{8\pi}u^{b}D_{a}v_{b}$ with $\mathfrak{K}=q^{ab}\mathfrak{K}_{ab}=q^{ab}D_{a}u_{b}$ are used in this context, where $v^{a}$ and $u^{a}$ are timelike and spacelike vector fields tangential and orthogonal to $\mathcal{B}$, $\mathfrak{K}_{ab}$ is the extrinsic curvature calculated with respect to $u^{a}$ and $D_{a}$ the covariant derviative at $\Omega_{t}$.
Sure enough, the form of the Hamiltonian depends on the choice of the time evolution vector field $t^{a}$. In principle, any given choice can reasonably be made for said vector field in absence of exact and/or asymptotic symmetries of spacetime. Yet, as for later purposes of this work, it will make sense to consider a null time-flow vector field of the form $t^{a}=\sqrt{2}Nk^{a}$; a choice that will only be legitimized in retrospect, in the further course of this section. For now, let it just be noted that the main motivation for this particular choice of $t^{a}$ comes from the theory of dynamical horizons, where a very similar choice is made for the time-flow vector field, but typically a specific ansatz for the lapse function of the geometry is made. In what follows, however, there is no need for such a specific choice of lapse, which is why no restrictions are placed on the latter.
Against this background, let it first be noted that the interior boundary Hamiltonian $H_{h}^{Boundary}$ vanishes as $\Xi=0$ applies. This can rather straightforwardly be concluded from the fact that, in view of the special choice made fo $t^{a}$, the boundary Hamiltonian density $\mathfrak{H}(\mathcal{N},\vec{\mathcal{N}})$ can be converted into
$\mathcal{\mathfrak{H}}=\frac{\sqrt{2q}N\Xi}{8\pi};$ (3)
see here [54] for further details. This form for the Hamiltonian density, as should be noted, proves consistent with the form of the quasilocal stress energy tensor calculated in [25] and [55], that is, to the corresponding expressions for the energy densities (see here equation (3.44)(3.44)( 3.44 ) on page 151515 of [25] and equation (36)(36)( 36 ) on page 161616 of [55], but note that slightly different null geometric settings and notational conventions are used).
To venture further, taking some other results of [53, 54] into account, let it be noted that the variation of the bulk part of the exterior Hamiltonian yields an integral expression of the form
$\mathfrak{L}_{t}H_{0}^{Bulk}=\underset{\Sigma_{t}}{\int}(\dot{N}\mathcal{H}+\mathcal{H}_{a}\dot{N}^{a}+\frac{N}{2}\mathcal{Q}_{ab}\dot{h}^{ab})\omega_{h}+\underset{\Omega_{t}}{\int}\Pi\omega_{q},$ (4)
where $\mathcal{Q}_{ab}=h_{a}^{\;c}h_{b}^{\;d}G_{cd}$ applies by definition, $\mathfrak{L}_{t}$ denotes the Lie derivative with respect to the time-flow vector field $t^{a}$ along $\Sigma_{t}$ and the occurring boundary integral is defined with respect to an integrand of the form $\Pi=\frac{N^{2}}{4\pi}(G_{ab}-8\pi T_{ab})k^{a}k^{b}$. Accordingly, this variation of the Hamiltonian vanishes on-shell, from which it can be concluded that, at the boundary of spacetime, the gravitational and matter bulk-to-boundary inflow terms occurring in (4)(4)( 4 ) cancel each other out. This, however, just implies the validity of the full Einstein equations
$G_{ab}=8\pi T_{ab}$ (5)
in the bulk and that of the associated scalar equation
$G_{ab}k^{a}k^{b}=8\pi T_{ab}k^{a}k^{b}$ (6)
at the boundary of spacetime. Of course, an analogous calculation can also be performed for the interior (horizon) part of the Hamiltonian $H_{h}$.
Meanwhile, as far as the exterior part of the Hamiltonian is concerned, a variation of the boundary part $H_{0}^{Boundary}$ yields a power functional $\mathcal{P}_{0}$ of the form
$\mathcal{P}_{0}=$ $\mathcal{L}_{t}H_{0}^{Boundary}=$ (7)
$=$ $\frac{1}{4\pi}\underset{\Omega_{t}}{\int}\left[\dot{\mathcal{N}}\mathfrak{h}+\mathfrak{h}_{a}\dot{\mathcal{N}}^{a}+\frac{\mathcal{N}}{2}\mathfrak{Q}_{ab}\dot{q}^{ab}\right]\omega_{q},$
given that $\mathcal{L}_{t}$ is the Lie derivative with respect to the time-flow vector field $t^{a}$ along $\Omega_{t}$. This result, as may be noted, is fully in line with known results from the literature [16] and has previosuly been rederived in [53, 54]. For the specific choice of $t^{a}=Nk^{a}$, as shown in [54], one finds
$\mathcal{P}_{0}=$ $\mathcal{L}_{t}H_{0}^{Boundary}=$ (8)
$=$ $\frac{1}{4\pi}\underset{\Omega_{t}}{\int}\left[N\mathcal{L}_{k}N\Xi+N^{2}(\varkappa\Xi-\sigma^{\prime}_{ab}\sigma^{\prime}{}^{ab}+\omega^{\prime}_{ab}\omega^{\prime}{}^{ab}-8\pi T_{ab}k^{a}k^{b})\right]\omega_{q},$
asssuming that $\varkappa=-(k\nabla)k_{a}l^{a}$ applies in this context. In the large sphere limit, this expression reduces to the form
$\mathcal{P}_{0,\infty}=\underset{r\rightarrow\infty}{\lim}\mathcal{L}_{t}H_{0}^{Boundary}$ $=$ (9)
=-= - $\frac{1}{4\pi}\underset{\mathcal{S}_{\infty}}{\int}\left[|\mathcal{\mathtt{N}}|^{2}+8\pi T_{kk}\right]d\Omega$
in the asymptotically flat context when standard fall-off conditions are assumed close to infinity (for further details, see here [48, 68]), where, as may be noted, the definitions $T_{kk}=\underset{r\rightarrow\infty}{\lim}r^{2}T_{ab}k^{a}k^{b}$ and $\sigma^{\prime}_{ab}\sigma^{\prime}{}^{ab}=|\sigma^{\prime}|^{2}=\frac{|\mathcal{\mathtt{N}}|^{2}}{r^{2}}$ have been used in this context. This makes it clear, however, that the expression obtained coincides exactly with the Bondi mass-loss formula discovered in [15], i.e. $\mathcal{P}_{0,\infty}\equiv\dot{M}_{B}$; see here [68] page 426426426 for comparison. As may be noted, this is fully consistent with a classic result of [23] (equation (0.1)(0.1)( 0.1 ) on page 333) and with the results of [55] (equation (57)(57)( 57 ) on page 101010), the subject of which is the integral law $\underset{\mathcal{S}_{\infty}}{\int}[\mathfrak{H}-\mathfrak{H}^{ref}]d^{2}x=\frac{1}{8\pi}\underset{\mathcal{S}_{\infty}}{\int}[\mathfrak{h}-\mathfrak{h}^{ref}]\omega_{q}=M_{B}$, which occurs for the choice $\mathcal{N}^{a}=0$, $\mathcal{N}=N=1$, where $M_{B}$ represents the Bondi mass and the reference term is chosen as that of flat Minokwski space. Consequently, by performing a variation of this expression with respect to the time-flow vector field chosen, one finds (since the variation of the boundary reference term is zero) integral law (9)(9)( 9 ) exactly reproduced.
Based on these findings, given two incoming and outgoing time-flow vector fields of the form $t_{\pm}^{a}=\sqrt{2}Nk_{\pm}^{a}$, the pair of quasilocal charges
$Q_{\pm}=\frac{1}{4\pi}\underset{\mathcal{I}^{\pm}}{\int}\left[|\mathcal{\mathtt{N}}_{\pm}|^{2}+8\pi T_{kk}^{\pm}\right]\varepsilon_{\pm q}$ (10)
can be specified by partial integration of (9)(9)( 9 ), where the definitions $T_{kk}^{\pm}:=\underset{r\rightarrow\infty}{\lim}r^{2}T_{ab}k_{\pm}^{a}k_{\pm}^{b}$ and $\varepsilon_{+q}=du\wedge\omega_{q}$ resp. $\varepsilon_{-q}=dv\wedge\omega_{q}$ have been used with $u$ and $v$ denote the associated retarded and advanced null time parameters, respectively.
1.3 Quasilocal Approach to Supertranslation Hair and Superrotation Charges, Antipodal Matching Conditions
In completely the same fashion as above, when making the choices $t_{+}^{a}=f\partial_{u}^{a}-\frac{1}{r}D^{A}f\partial_{A}^{a}+\frac{1}{2}D^{2}f\partial_{r}^{a}=fk_{+}^{a}+V_{+}^{a}$ with $\partial_{u}f=\partial_{r}f=0$ and $t_{-}^{a}=f\partial_{v}^{a}-\frac{1}{r}D^{A}f\partial_{A}^{a}+\frac{1}{2}D^{2}f\partial_{r}^{a}=fk_{-}^{a}+V_{-}^{a}$ with $\partial_{v}f=\partial_{r}f=0$, the quasilocal charges
$Q_{f}^{\pm}=\frac{1}{4\pi}\underset{\mathcal{I}^{\pm}}{\int}\left[|\mathcal{\mathtt{N}}_{\pm}|^{2}-\frac{1}{4}D_{a}D_{b}\mathtt{N}_{\pm}^{ab}+8\pi T_{kk}^{\pm}\right]f\varepsilon_{\pm q}$ (11)
can be deduced from (8)(8)( 8 ) by taking the large sphere limit in Bondi coordinates close to null infinity, where some important details of the calculation are given in Appendix $A$ of this work. To convince oneself about the consistency of this result, one may take another result of [23] into account (equation (4.5)(4.5)( 4.5 ) on page 999), which shows that for a time-flow vector fields meeting the condition $t_{+}^{a}=f\partial_{u}^{a}|_{\mathcal{I}^{+}}$ and $t_{-}^{a}=f\partial_{v}^{a}|_{\mathcal{I}^{-}}$ as well $\mathcal{N}^{a}=0$ at the spatial boundary $\mathcal{B}$ of spacetime it follows at leading order that
$Q_{f}^{\pm}=\underset{r\rightarrow\infty}{\lim}H_{0}^{Boundary}=\frac{1}{4\pi}\underset{\mathcal{S}_{\infty}}{\int}fm_{B}^{\pm}\omega_{q},$ (12)
for $H_{0}^{Boundary}=\frac{1}{8\pi}\underset{\mathcal{S}_{\infty}}{\int}[\mathfrak{h}-\mathfrak{h}^{ref}]\omega_{q}$, $f(x^{2},x^{3})=\underset{r\rightarrow\infty}{\lim}N(r,x^{2},x^{3})$, given that $m_{B}(u,x^{2},x^{3})$ resp. $m_{B}(v,x^{2},x^{3})$ is the Bondi mass aspect of the geometry, thus giving an expression that corresponds to Geroch’s supermomentum [23, 37] in case that the boundary reference term is chosen appropriately. Yet, this allows one to conclude that
$Q_{f}^{\pm}=\underset{r\rightarrow\infty}{\lim}H_{0}^{Boundary}=\frac{1}{4\pi}\underset{\mathcal{I}^{\pm}}{\int}f\dot{m}_{B}^{\pm}\varepsilon_{\pm q},$ (13)
giving an expression that immediately leads to (11)(11)( 11 ) by virtue of the fact that $\partial_{u}m_{B}^{+}=|\mathcal{\mathtt{N}}_{+}|^{2}+\frac{1}{4}D_{a}D_{b}\mathtt{N}_{+}^{ab}+8\pi\mathfrak{T}_{uu}$ resp. $\partial_{v}m_{B}^{-}=|\mathcal{\mathtt{N}}_{-}|^{2}+\frac{1}{4}D_{a}D_{b}\mathtt{N_{-}}^{ab}+8\pi\mathfrak{T}_{vv}$ applies in Bondi coordinates close to null infinity if the fall-off conditions presented in Appendix A of this work are taken into account. That said, it becomes clear that the quasilocal charges derived are manifestly identical with the soft supertranslation charges derived in [32, 46], thus making it clear that the authors’ results can be reproduced also within the Brown-York quasilocal framework used. As may also be noted, the deduced quasilocal charges also turn out to be identical in the asymptotic regime to those constructed in [31], even though a different quasilocal approach was adopted in the mentioned work.
For all these agreements, however, it is important to note that no considerations regarding bulk symmetries or conserved charges on the horizon were made. Only the asymptotic behavior of quasilocal Brown-York charges resp. their time derivatives was investigated, without imposing any specific constraints on the spacetime geometry. In that sense, the results obtained prove to be both quite general and fully consistent with previous results from the literature.
[…]
$\underset{\Omega_{t}}{\int}\mathfrak{h}_{a}\mathcal{N}^{a}\omega_{q}=\frac{1}{8\pi}\underset{\Omega_{t}}{\int}[\Omega\Omega_{a}\varphi^{a}+\Omega(\varphi D)\ln\xi]\omega_{q}=J_{\mathcal{S}_{t}}^{\Omega\varphi}+\frac{1}{8\pi}\underset{\Omega_{t}}{\int}\Omega(\varphi D)\ln\xi\omega_{q}.$ (14)
Taking then the large sphere limit and assuming that the relation
$\Omega\Omega_{a}\varphi^{a}=\mathcal{\mathtt{N}}_{a}Y^{a}$ (15)
admits a solution close to null infnity (with this being a clearly justified assumption as $\Omega\varphi^{a}$ can in principle be chosen arbitrarily), a suitable choice for the gauge term occurring in (14)(14)( 14 ) allows the further pair of charges
$Q_{Y}^{+}=\frac{1}{8\pi}\underset{\mathcal{I}_{-}^{+}}{\int}\mathcal{\mathtt{N}}_{a}Y^{a}\omega_{q},\;\frac{1}{8\pi}\underset{\mathcal{I}_{+}^{-}}{\int}\mathcal{\mathtt{N}}_{a}Y^{a}\omega_{q}=Q_{Y}^{-},$ (16)
to be defined, which coincide again exactly with the expressions given in [32, 46]. Here, as may be noted, the vector field $\mathcal{\mathtt{N}}_{a}$ is nothing but the the angular momentum aspect of spacetime and $Y^{a}$ is a conformal Killing vector field on the two-sphere given by $Y^{a}=D^{a}\chi+\varepsilon^{ab}D_{b}\kappa$ with $\chi$ and $\kappa$ representing solutions of the differential relation $(D^{2}+2)\chi=(D^{2}+2)\kappa=0$ for $D^{2}:=D_{a}D^{a}.$
To summarize, the preceding arguments show that a variation of the Brown-York-Hamiltonian (2)(2)( 2 ) yields on-shell expressions for supertranslation charges (11)(11)( 11 ) and superrotation charges (16)(16)( 16 ) in null coordinates in the Bondi gauge, both of which agree exactly with known results from the literature.
As pointed out in [46], the validity of the antipodal matching conditions
$Q_{f,Y}^{+}=Q_{f,Y}^{-}$ (17)
ensures that the gravitational scattering problem is well-defined, while additionally implying that an infinite number of charges are conserved in relativsitic scattering processes. Such a condition therefore constitutes a natural starting point for finding a possible resolution to the information loss paradox.
As far as the latter is concerned, however, the situation is somewhat complicated in that in stationary black hole spacetimes $\mathcal{I}^{-}$ happens to be a Cauchy hypersurface, whereas $\mathcal{I}^{+}$ fails to be one. Still, the hypersurface $\mathcal{I}^{+}\cup\mathcal{H}^{+}$ constitutes a Cauchy surface, so that it makes sense - in the case of massless fields - to infer that $Q_{f}^{-}=Q_{f}^{\mathcal{I}^{-}}$ and $Q_{f}^{+}=Q_{f}^{\mathcal{I}^{+}}+Q_{f}^{\mathcal{H}^{+}}$, thus leading one to conclude that
$Q_{f}^{\mathcal{H}^{+}}=Q_{f}^{\mathcal{I}^{-}}-Q_{f}^{\mathcal{I}^{+}}.$ (18)
The resultant charge $Q_{f}^{\mathcal{H}^{+}}$ represents a major obstacle to solving Hawking’s paradox, for its existence is linked to an increase in entropy, and thus a loss of information, at the quantum level.
The idea expressed in [46] to overcome this problem is now to consider a supertranslated black hole spacetime resulting from ’throwing in’ an asymmetric shock wave, and to explicitly determine the form of the resultant charge (18)(18)( 18 ); a thought-provoking approach that should clearly prove consistent with the results of the present section. Notwithstanding, said approach will not be pursued further below.
Instead, the idea will be pursued of modelling - in full agreement with the assumptions made in this section - the black hole horizon as a dynamical horizon in the sense of Ashtekar et al. (resp. future trapping in the sense of [48]). This changes the entire setting, as now the black hole horizon is no longer modelled for all times as a quasi-stationary hypersurface through which neither matter nor radiation can escape. On the contrary, the horizon is now modeled as a dynamical non-teleological geometric object that allows Hawking radiation to escape to infinity. In combination with specific fall-off conditions on the field modes of the scalar field, which are derived in Appendix B by means of the model of a scalar field in a Vaidya sandwich spacetime, this change of setup proves sufficient to ensure that an information loss paradox in the sense of Hawking should not occur. The main reason for this is that in the given setting the black hole horizon is a dynamical horizon, which permits radiation modes to escape and carry quasilocal energy, and hence information, through the black hole horizon to infinity, thus allowing all information encoded in the emitted radiation in principle to be recovered at null infinity. Before going into details at this point, however, this section will close by demonstrating that the results obtained so far can be generalized in several respects by taking the results of [54] into account, the latter of which can be used to construct quasilocal charges of a more general form than (10)(10)( 10 ) and (11)(11)( 11 ).
1.4 Generalized Results: Mass and Radiation Transfer in Bounded and Unbounded Gravitational Fields
As a basis for constructing quasilocal charges that generalize those depicted in (10)(10)( 10 ) and (11)(11)( 11 ), a generic ansatz of the form $t^{a}=\sqrt{2}Nk^{a}+V^{a}$ for the time evolution vector of spacetime will be made in the following, where $V^{a}=-Ns^{a}+N^{a}=(O-N)s^{a}+\mathcal{N}^{a}$ has to be satisfied for the sake of consistency.
Knowing that a time-flow vector field for spacetimes without time-translation symmetry can arguably almost arbitrarily be chosen, the question arises as to which ansatz for the shift vector field $V^{a}$ should ideally be made. One way to approach this question is to first consider a black hole spacetime $(\mathcal{M},g)$ with time translation symmetry and axial symmetry, according to which the time-flow vector field can simply be chosen to be the Killing vector field of the geometry (that is, a linear combination of temporal and angular Killing vector fields). In such a geometric setting, it certainly makes sense to make an ansatz of the form $V^{a}=\Omega\varphi^{a}$ close to the Killing horizon of spacetime, where $\Omega$ and $\varphi^{a}$ constutite the angular velocity of the black hole and the angular Killing vector field of the geometry. Admittedly, the resulting quasilocal charge can then also be meaningfully defined at null infinity. This, however, should apparantly work out in the more general case of a dynamical black hole spacetime, which is intriguing not least because, as shown in [54], the ansatz made allows the derivation of the Ashtekar-Krishnan version of the first law of black hole mechanics (with respect to a part $\Delta\mathcal{H}$ of the dynamical horizon $\mathcal{H}$) [6].
That said, given the above vector field, the boundary Hamiltonian density takes the form
$\mathcal{\mathfrak{H}}=\frac{\sqrt{q}}{8\pi}\left[\sqrt{2}N\Xi-\varGamma_{V}\right],$ (19)
by virtue that the definition $\varGamma_{V}=(K_{ab}-Kh_{ab})s^{a}V^{b}+\lambda(\mathcal{N}\mathcal{D})\eta$ is used in the present context. By a temporal variation of the corresponding boundary term $H_{0}^{Boundary}$, one obtains from the above - taking into account that $s^{a}=\frac{1}{\sqrt{2}}(k^{a}-l^{a})$ and $D_{a}s^{a}=q^{ab}\nabla_{a}s_{b}=\frac{1}{\sqrt{2}}(\Xi-\Theta)$ applies by definition - the generalized power functional
$\mathscr{P}_{0}=\mathcal{L}_{t}H_{0}^{Boundary}=$ $\mathcal{P}_{0}+\mathcal{L}_{V}\underset{\Omega_{t}}{\int}\mathcal{\mathfrak{H}}\omega_{q}=$ (20)
$=\mathcal{P}_{0}+\sqrt{2}\underset{\Omega_{t}}{\int}(O-N)[\mathcal{L}_{k-l}\mathcal{\mathfrak{H}}$ $+(\Xi-\Theta)\mathcal{\mathfrak{H}}]\omega_{q}+\mathcal{L}_{\mathcal{N}}\underset{\Omega_{t}}{\int}\mathcal{\mathfrak{H}}\omega_{q}$
where $\mathcal{P}_{0}$ is given by (7)(7)( 7 ). Yet, taking the fact into account that an integral over the boundary of a boundary is zero, one finds that
$\mathcal{L}_{\mathcal{N}}\underset{\Omega_{t}}{\int}\mathcal{\mathfrak{H}}\omega_{q}=\underset{\Omega_{t}}{\int}D_{a}(\mathcal{\mathfrak{H}}\mathcal{N}^{a})\omega_{q}=0$ (21)
applies to the corresponding integral expression depicted in (20)(20)( 20 ). Accordingly, one is thus led to onclude that relation (20)(20)( 20 ) can be re-written in the form
$\mathscr{P}_{0}=\mathcal{P}_{0}+\mathfrak{P}_{0},$ (22)
provided that the definition
$\mathfrak{P}_{0}=\sqrt{2}\underset{\Omega_{t}}{\int}(O-N)[\mathcal{L}_{k-l}\mathcal{\mathfrak{H}}+(\Xi-\Theta)\mathcal{\mathfrak{H}}]\omega_{q}$ (23)
is used in the present context. By performing the large sphere limit, then, relation (22)(22)( 22 ) can be further converted by taking (9)(9)( 9 ) into account, thereby yielding the result
$\mathscr{P}_{0,\infty}=\dot{M}_{B}+\mathfrak{P}_{0,\infty},$ (24)
where $M_{B}$ is the Bondi-Sachs mass. Consequently, one is lead to conclude that the occurring expression $\mathfrak{P}_{0,\infty}=\underset{r\rightarrow\infty}{\lim}\mathfrak{P}_{0}$ with $\mathfrak{P}_{0}$ being given by (23)(23)( 23 ) - gives rise to quasilocal corrections to the Bondi mass loss formula, which, as first shown in [54], can be expected to be zero if and only if the function $O(x)$ and the boundary shift vector field $\mathcal{N}^{a}(x)$ are chosen appropriately. In general, however, the derived quasilocal corrections will be different from zero. Their existence thus reveals that in generic geometric settings, there should, according to the Brown-York quasilocal formalism, be exceptions from the generally accepted rule: The mass of a system is constant if and only if there is no news.
The derived corrections should be taken into account, since by integrating (24)(24)( 24 ) they allow the definition of the class of quasilocal charges, which contains the class of charges given by (11)(11)( 11 ) as a special case. Also, as may be noted, these corrections play a role in the variation of the boundary part of the horizon Hamiltonian $H_{h}^{Boundary}$ (recall that the variation of the bulk part $H_{h}^{Bulk}$ vanishes on-shell). A possible ansatz for the time-flow vector field leading to such a horizon Hamiltonian is $t^{a}=\sqrt{2}Nk^{a}+V^{a}$ with $V^{a}=\mathcal{O}s^{a}+2\Omega\varphi^{a}$ and $O=N+\mathcal{O}$ such that $\mathcal{O}=\frac{\varkappa}{4\pi(K-K_{ab}s^{a}s^{b})}$, thereby making it clear that $\varkappa$ coincides with the surface gravity of the black hole, where one has thus $-\varkappa=\gamma+\bar{\gamma}$ in spin-coefficients. The boundary part of the Hamiltonian then takes the form
$H_{h}^{Boundary}=\underset{\mathcal{S}_{t}}{\int}\mathfrak{H}d^{2}x=\underset{\mathcal{S}_{t}}{\int}[\sqrt{2}N\Xi+\frac{\varkappa}{8\pi}+2\Omega\Omega_{a}\varphi^{a}]\omega_{q}+\varDelta$ (25)
where the relations $\mathcal{N}\mathfrak{K}=Nk-(K_{ab}-Kh_{ab})N^{a}s^{b}-\lambda(\mathcal{N}\mathcal{D})\eta-(\mathcal{N}\mathcal{D})v_{a}u^{a}$ and $\mathfrak{h}_{a}\mathcal{N}^{a}=2\Omega[\Omega_{a}\varphi^{a}-(\phi D)\ln\xi]$ have been used to obtain this particular form of the result.
As may be noted, the $\varDelta$-term occurring in (25)(25)( 25 ) is a gauge-dependent term that can be eliminated by a suitable choice for $\xi(x)$. Thus, in case that the latter is assumed to be so and, moreover, $(\mathcal{M},g)$ is assumed to exhibit an axisymmetric vacuum black hole four-geometry that is invariant under time translations, where $\{\mathcal{S}_{t}\}$ shall be the folia of a non-expanding Killing horizon $\mathcal{H}^{+}$ that forms the interior boundary of spacetime and $\mathcal{E}$ denotes the Komar energy evaluated at the horizon, equation (25)(25)( 25 ) can be re-written in the form
$H_{h}^{Boundary}=\frac{\varkappa}{4\pi}\mathcal{A}+2\Omega\mathcal{J}=\mathcal{E},$ (26)
where it has been used that $\Xi=0$ as well as $\varkappa=const.$ and $\Omega=const.$ applies locally at the black hole horizon. Of course, $\mathcal{A}$ and $\mathcal{J}$ denote here the area and angular momentum of the black hole, respectively.
Accordingly, it can be concluded that the Brown-York Hamiltonian agrees with the Komar energy in the given case except for one eliminable gauge term, thereby allowing Smarr’s formula and the first law of black hole mechanics to be reproduced (given that the black hole is uncharged). In consequence, however, a variation (not necessarily a temporal variation) of said boundary term yields the well-known result
$\delta\mathcal{E}=\frac{\varkappa}{8\pi}\delta\mathcal{A}+\Omega\delta\mathcal{J},$ (27)
provided that $N\delta\Xi$ is locally zero at the sections of the horizon (as would be the case for a temporal variation of (26)(26)( 26 )).
As may be noted, an analogous scenario occurs in the electrovac case resp. even when the notion of the Killing Horizon is replaced by a more general non-expanding null horizon (i.e. either a weakly isolated or isolated horizon in the sense of [7]), leading to the generalized first law of black hole mechanics derived in [2, 3]. In the latter case, as explained in more detail in the upcoming section, it then would have to be expected that corrections of the form (24)(24)( 24 ) occur at infinity, at least provided that the weakly isolated or isolated horizon results as a special case of a dynamical horizon.
As described at the beginning of this section, such a dynamical horizon constitutes the inner boundary of spacetime in the introduced geometric setting and therefore forms a natural part of the presented model, while in contrast no such horizon was considered in [46]. Yet, as will now be shown, the existence of such a horizon makes a significant difference in modeling black hole formation and evaporation processes. For according to such a scenario, no information loss paradox occurs, since in the phases in which spacetime is nonstationary and in which the black hole horizon is dynamical, radiation modes can carry away quasilocal energy and thus information through the black hole horizon to null infinity, thereby causing backreactions on the black hole singularity until the latter vanishes completely. This is to be explained in more detail below.
2 Brown-York charges and Black Hole Soft Hair: Classical and Quantum fields
The arguments presented so far have been entirely classical. This is now to change. To this end, the framework introduced in the previous section will be revisited below within the semiclassical framework of Euclidean gravity. In doing so, it is shown that quasilocal charges of the form (10)(10)( 10 ), (11)(11)( 11 ) and (16)(16)( 16 ) can be constructed analogously in the semiclassical approximation to quantum gravity, including the important special case of a quantized massless scalar field in a nonstationary black hole spacetime. Furthermore, it is shown that the antipodal matching conditions introduced in the previous section can be satisfied if the black hole horizon is assumed to be given by a dynamical horizon resp. a future trapping horizon in the sense of [1, 49]. To satisfy these matching conditions, constraints are imposed on the mode function of the quantized massless scalar field, which are justified in Appendix B using the simplified example of a scalar field in a Vaidya background spacetime with a time-dependent monotonically decreasing mass that becomes zero at the very end of the evaporation process. Based on the results obtained, it is eventually concluded that particle creation effects in nonstationary black hole spacetimes with dynamical horizons should not cause information loss at the quantum level.
2.1 Nonstationary Black Holes, Particle Creation Effects and Information Loss in a Nutshell
To generalize the results obtained in the previous section and to include quantum effects in the physical framework presented so far, a brief overview of particle creation effects in curved spacetimes is given in this subsection. To this end, the geometric setting considered so far is modified in that a sandwich spacetime is now taken into account, which describes a black hole formed by the collapse of a massless scalar field that evaporates after a certain time. That is to say, a spacetime $(M,g)$ is considered whose metric can be treated essentially classically; at least as long as quantum backreaction effects are not taken into account, which shall be assumed for the time being. This spacetime contains three different regions: $i)$ an initial flat region, $ii)$ a dynamical curved black hole region, and eventually again $iii)$ a final flat region. Concerning the considered matter fields, the focus will be placed - similarly to Hawking’s original work [43] on the subject - on the case of a massless, quantized scalar field $\varphi(x)$, which constitutes a solution of the covariant wave equation
$\boxempty\varphi=\nabla_{a}\nabla^{a}\varphi=0.$ (28)
As for any field in a generic curved geometric background, solutions of said equation cannot be decomposed into its positive and negative frequency parts, since positive and negative frequencies have no invariant meaning in a curved spacetime when the latter lacks time translation symmetry. Yet, since the spacetime model considered contains an initial flat region, one can find the decomposition
$\varphi=\overset{\infty}{\underset{k=1}{\sum}}[f_{k}^{-}a_{k}^{-\dagger}+\bar{f}_{k}^{-}a_{k}^{-}]$ (29)
of $\varphi(x)$ such that the purely ingoing mode functions $\{f_{k}^{-}(x)\}$ and $\{f_{k}^{-\ast}(x)\}$ contain only positive frequencies with respect to the initial retarded Minkowskian time coordinate. Here, $a_{k}^{-}$ and $a_{k}^{-\dagger}$ are latter operators satisfying the standard canonical commutation relations for massless bosonic fields.
The above decomposition of $\varphi(x)$ is of course not unique. Still, a decomposition of $\varphi(x)$ with respect to the final retarded Minkowskian time coordinate and associated purely outgoing field modes $\{f_{k}^{+}(x)\}$ and $\{f_{k}^{+\ast}(x)\}$ generally cannot be performed due to the fact that $\mathcal{I}^{+}$, quite in contrast to $\mathcal{I}^{-}$, fails to be a Cauchy hypersurface. One way out of this dilemma is to carry out the decomposition of $\varphi(x)$ with respect to $\mathcal{I}^{+}\cup\mathcal{H}$, where $\mathcal{H}$ is the black hole horizon. The reason is that $\mathcal{I}^{+}\cup\mathcal{H}$ is actually a Cauchy horizon. Given this particular choice, the decomposition reads
$\varphi=\overset{\infty}{\underset{k=1}{\sum}}[f_{k}^{+}a_{k}^{+\dagger}+\bar{f}_{k}^{+}a_{k}^{+}+g_{k}b_{k}^{\dagger}+\bar{g}_{k}b_{k}],$ (30)
where $\{g_{k}(x)\}$ and $\{g_{k}^{\ast}(x)\}$ are horizon modes and $a_{k}^{+\dagger}$, $a_{k}^{+}$ and $b_{k}^{\dagger}$, $b_{k}$ are creation and annihilation operators satisying the canonical commutation relations
$[b_{k},b_{l}]=[b_{k},a_{l}^{\pm}]=[b_{k},a_{l}^{\pm\dagger}]=[a_{k}^{\pm},a_{l}^{\pm}]=0,\;[a_{k}^{\pm},a_{l}^{\pm\dagger}]=[b_{k},b_{l}^{\dagger}]=\delta_{kl}.$ (31)
Given that spacetime is flat at early and late times, it can thus be concluded that the pairs of annihilation operators $a_{k}^{\pm}$ give rise to initial and final asymptotic vacuum states $|0\rangle_{\pm}$, which both states are solutions of the equations $a_{k}^{\pm}|0\rangle_{\pm}=0$. Since the condition of asymptotic completeness [40, 47] is however clearly not fulfilled under the given geometric circumstances (as the model considered lacks a unique Poincaré invariant vacuum state), these asymptotic vacuum states must be inherently different from each other. Also the form of the final mode functions $\{f_{k}^{+}(x)\}$ will differ from that of the initial mode functions $\{f_{k}^{-}(x)\}$. Yet, since the latter form a complete basis for solutions of the wave equation (28)(28)( 28 ) and massless fields are completely determined by their initial data at past null infinity, the mode functions and ladder operators in the future asymptotic region are related with those in the past asymptotic region by the Bogoliubov transformations
$f_{k}^{+}=\overset{\infty}{\underset{m=1}{\sum}}(\alpha_{km}f_{m}^{-}+\beta_{km}\bar{f}_{m}^{-}),\;g_{k}=\overset{\infty}{\underset{m=1}{\sum}}(\gamma_{km}f_{m}^{-}+\eta_{km}\bar{f}_{m}^{-})$ (32)
and
$a_{k}^{+}=\overset{\infty}{\underset{m=1}{\sum}}(\alpha_{km}^{*}a_{m}^{-}-\beta_{km}^{*}a_{m}^{-\dagger}),\;b_{k}=\overset{\infty}{\underset{m=1}{\sum}}(\gamma_{km}^{*}a_{m}^{-}-\eta_{km}^{*}a_{m}^{-\dagger}),$ (33)
where the onditions $\overset{\infty}{\underset{m=1}{\sum}}(\alpha_{km}\alpha_{mj}^{*}-\beta_{km}\beta_{mj}^{*})=\delta_{kj}$ and $\overset{\infty}{\underset{m=1}{\sum}}(\alpha_{km}\beta_{mj}-\beta_{km}\alpha_{mj})=0$ as well as $\overset{\infty}{\underset{m=1}{\sum}}(\gamma_{km}\gamma_{mj}^{*}-\eta_{km}\eta_{mj}^{*})=\delta_{kj}$ and $\overset{\infty}{\underset{m=1}{\sum}}(\gamma_{km}\eta_{mj}-\eta_{km}\gamma_{mj})=0$ need to be satisfied for the sake of consistency. Solving the relation $a_{k}^{+}|0\rangle_{+}=\overset{\infty}{\underset{m=1}{\sum}}(\alpha_{km}^{*}a_{m}^{-}-\beta_{km}^{*}a_{m}^{-\dagger})|0\rangle_{+}=0$ then yields
$|0\rangle_{+}=Nexp(-\frac{1}{2}\overset{\infty}{\underset{m,n=1}{\sum}}\chi_{mn}a_{m}^{-\dagger}a_{n}^{-\dagger})|0\rangle_{-}$ (34)
for some normalization factor $N$ with $\chi=\frac{1}{2}(\alpha^{-1}\beta+(\alpha^{-1}\beta)^{T})$, thereby showing that the asymptotic vacuum ground states of the field do not coincide.
[…]
$N_{k}={}_{-}\langle 0|a_{k}^{+\dagger}a_{k}^{+}|0\rangle_{-}=\overset{\infty}{\underset{l=1}{\sum}}|\beta_{kl}|^{2}=\overset{\infty}{\underset{l=1}{\sum}}\frac{\Gamma_{kl}}{e^{\frac{2\pi\omega_{k}}{\kappa}}-1}.$ (35)
The semiclassical arguments used thus clearly indicate that the final quantum state of the field should be a mixed thermal state with Planck spectrum of the form (35)(35)( 35 ) after the black hole has evaporated completely, where $\kappa$ is the surface gravity of the black hole. One therefore is led to conclude that that a given pure initial quantum state $\rho_{in}\equiv|\Psi\rangle\langle\Psi|$, where $|\Psi\rangle$ is some Fock state constructed from $|0\rangle_{-}$, inevitably evolves into a mixed state $\rho_{out}$ in the course of black hole evaporation. The latter leads to an increase in entropy and thus to a loss of quantum information; something that seems to conflict with the standard rules of quantum field theory in flat Minkowski space.
That said, the exact way in which this loss of quantum information manifests itself with respect to the quasilocal charges constructed in the previous section will now be eludicated in the upcoming subsection of this work, where charges of the form (10)(10)( 10 ), (11)(11)( 11 ) and (16)(16)( 16 ) are constructed in the framework of semiclassical Euclidean quantum gravity. In doing so, as will be shown, an essential step is to take into account those quantum backreaction effects that have remained unconsidered in the present subsection.
2.2 Boundary Charges and Quantum Backreaction
So far, quantum backreaction effects of matter on the geometry of spacetime have been neglected. In the short term, this may well be justified for black holes in quasi-stationary equilibrium. Since the continuous emission of Hawking radiation due to particle creation effects clearly leads to a change in the geometry of the black hole over time, however, these effects cannot simply be ignored in the long run. On the contrary, these quantum backreaction effects must clearly be taken into account in the definition of quasilocal Brown-York charges; now to be generalized in the semiclassical framework of Euclidean quantum gravity.
As a basis for the latter, it may be noted that the stress-energy operator corresponding to a quantized massless scalar field $\varphi(x)$ of the form (28)(28)( 28 ) reads
$T_{ab}=\nabla_{a}\varphi\nabla_{b}\varphi,$ (36)
meaning that one actually considers a distributional bitensor field $T_{ab}(x,x^{\prime})$ taking values at two different spacetime points $x$ and $x^{\prime}$, which results from applying a bidifferential operator on the bilinear product of operator-valued distributions (the field operators). To obtain a well-defined tensor distribution from this expression that is solely defined at a single point, it has proven convenient to calculate the renormalized expectation value $\langle T_{ab}\rangle_{ren}(x)\equiv\langle T_{ab}\rangle_{\omega,ren}(x)$ of $T_{ab}(x,x^{\prime})$ in some (algebraic) state $\omega$, where the same state can be used as a basis for a GNS construction of the corresponding Hilbert space. The calculation of $\langle T_{ab}\rangle_{ren}(x)$, however, usually proves to be a subtle, technically involved issue, which is essentially due to the fact that the point limit $x\rightarrow x^{\prime}$ of $T_{ab}(x,x^{\prime})$ is singular. Notwithstanding this, said expectation value can be calculated by different techniques known from the literature (see here, for example, [14, 35, 65, 76] and references therein for further information), which allow the divergent parts of $\langle T_{ab}\rangle(x)$ to be removed, thereby giving a suitably regularized version of this expectation value a mathematically well defined meaning. Whereas some of these techniques focus primarily on the renormalization of coupling constants in the gravitational action, others focus more on controlling the short distance behavior of the two-point distribution $\langle\varphi(x)\varphi(x^{\prime})\rangle$ in order to calculate $\langle T_{ab}\rangle_{ren}(x)$ from $T_{ab}(x,x^{\prime})$ in the point limit $x\rightarrow x^{\prime}$. These latter so-called point-splitting methods [27, 28, 30, 35, 36, 76] additionally reveal the fact that $\langle T_{ab}\rangle_{ren}(x)$ can only be reasonably calculated in a suitable class of states, which, as is generally agreed, is the class of (quasi-free) Hadamard states. These states not only allow the expectation value of the stress-energy tensor to be calculated in rigorous fashion (e.g. by using the Hadamard renormalization approach), but also show - by Wald’s theorem [75, 76] - that $\langle T_{ab}\rangle_{ren}(x)$ is only uniquely defined up to the ambiguity of adding local curvature terms.
This all plays a role for the generalization of the quasilocal Brown-York charges derived in the previous section. The reason for this is that knowledge of the expectation value $\langle T_{ab}\rangle_{ren}$ is crucial for the description of quantum backreaction effects on the geometry of spacetime in the context of the semiclassical Euclidean quantum gravity. This, in turn, is due to the fact that the central role in such a description is played by the semiclassical Einstein equations
$G_{\;b}^{a}=8\pi\mathcal{T}_{\;b}^{a}=8\pi\langle T_{\;b}^{a}\rangle_{ren}-\varTheta_{\;b}^{a},$ (37)
which differ by higher order corrections from the classical Einstein equations (5)(5)( 5 ). These higher order corrections are all encoded in the object $\mathcal{T}{}_{ab}=\langle T_{ab}\rangle_{ren}-\frac{1}{8\pi}\varTheta_{ab}$; respectively, to be more precise, in the object
$\varTheta_{\;b}^{a}=c_{1}H_{1\;b}^{a}+c_{2}H_{2\;b}^{a}+c_{3}H_{3\;b}^{a},$ (38)
which is given with respect to the renormalization constants $c_{1}$, $c_{2}$ and $c_{3}$ and the tensor fields $H_{1\;b}^{a}$, $H_{2\;b}^{a}$ and $H_{3\;b}^{a}$
$H_{1\;b}^{a}$ $=2\nabla^{a}\nabla_{b}R-2\boxempty R\delta_{\;b}^{a}+2RR_{\;b}^{a}-\frac{1}{2}R^{2}\delta_{\;b}^{a},$ (39)
$H_{2\;b}^{a}$ $=2\nabla_{b}\nabla^{c}R_{\;c}^{a}-\boxempty G_{\;b}^{a}+2R_{\;c}^{a}R_{\;b}^{c}-\frac{1}{2}\delta_{\;b}^{a}R_{\;d}^{c}R_{\;c}^{d},$
$H_{\epsilon 3\;b}^{a}$ $=-\frac{1}{2}\delta_{\;b}^{a}R_{cdef}R^{cdef}+2R_{bcde}R^{acde}-4\boxempty R_{\;b}^{a}+2\nabla^{a}\nabla_{b}R$
$-4R_{\;c}^{a}R_{\;b}^{c}+4R^{cd}R_{\;cbd}^{a};$
see here, for instance, [14] for further information.
To set up (37)(37)( 37 ), it must therefore be ensured that $\langle T_{ab}\rangle_{ren}(x)$ and thus $\mathcal{T}_{\;b}^{a}$ can actually be calculated. If the latter is the case, the obvious idea of how to transition from the classical Brown-York charges derived in the previous section to the corresponding quantum charges is to make the replacement $T_{\;b}^{a}\rightarrow\mathcal{T}_{\;b}^{a}$ in relations (410)(4-10)( 4 - 10 ). This yields, among other things, the result
$\mathcal{L}_{t}$ $H_{h}^{Boundary}=$ (40)
$=\frac{1}{4\pi}\underset{\mathcal{S}_{t}}{\int}$ $\left[N\mathcal{L}_{k}N\Xi+N^{2}[\varkappa\Xi-\sigma^{\prime}_{ab}\sigma^{\prime}{}^{ab}+\omega^{\prime}_{ab}\omega^{\prime}{}^{ab}-8\pi\mathcal{T}{}_{ab}k^{a}k^{b}]\right]\omega_{q},$
with $\mathcal{S}_{t}\subset\mathcal{H}$ and thus
$\mathcal{L}_{t}H_{h}^{Boundary}$ $=$ (41)
=-= - $\frac{1}{4\pi}\underset{\mathcal{S}_{\infty}}{\int}\left[|\mathcal{\mathtt{N}}|^{2}+8\pi\mathcal{T}{}_{kk})\right]d\Omega,$
in the large sphere limit, where the definition $\mathcal{T}{}_{kk}:=\underset{r\rightarrow\infty}{\lim}[r^{2}(\langle T_{ab}\rangle_{ren}k^{a}k^{b}-\varTheta_{ab}k^{a}k^{b})]$ has been used.
So as to be able to consider said charges also within the semiclassical regime of Euclidean quantum gravity for the given geometric model of a collapsing and evaporating black hole spacetime, it shall be assumed in the following that the renormalized expectation value of the stress-energy tensor can actually be calculated at early and late times (i.e. near past and future null infinity) from the asymptotic vacuum states $|0\rangle_{\pm}$. That this is indeed possible is by no means obvious, as the geometric model considered is dynamical and thus lacks temporal and spatial translation symmetry as well as a preferred Poincaré invariant vacuum state. Notwithstanding that, it may however be possible to single out within the flat asymptotic regions of spacetime the preferred asymptotic vacuum states $|0\rangle_{\pm}$ as well as associated preferred Hadamard states $\omega^{\pm}$ relative to which the pairs $\langle T_{ab}^{\pm}\rangle_{ren}:=\langle T_{ab}\rangle_{\omega^{\pm},ren}$ and $\mathcal{T}_{ab}^{\pm}:=\langle T_{ab}^{\pm}\rangle_{ren}-\frac{1}{8\pi}\varTheta_{ab}$ of tensor distributions can be calculated (whereby, of course, one and the same renormalization scheme must be used).
Although the assumptions made are, as can be seen, quite strong, it should be noted that they are both legitimate and in some ways even substantial. For without a corresponding selection of asymptotic states with a reasonably decent physical behavior, it would be extremely hard to even posit Hawking’s information paradox, let alone discuss quantum backreaction effects and their physical consequences for the evolution and ultimate fate of an evaporating black hole spacetime. The assertions made therefore seem justified at least to the extent that they allow particle creation effects and any alleged loss of quantum information resulting from them to be discussed.
[…]
±$\underset{\mathcal{I}^{\pm}}{\int}N^{2}\dot{\Xi}\varepsilon_{\pm q}=\underset{\mathcal{I}^{\pm}}{\int}N^{2}[\kappa\Xi-\Xi^{2}-\sigma^{\prime}_{ab}\sigma^{\prime}{}^{ab}+\omega^{\prime}_{ab}\omega^{\prime}{}^{ab}-8\pi G{}_{k_{\pm}k_{\pm}}]\varepsilon_{\pm q},$ (42)
provided that the definition $G_{k_{\pm}k_{\pm}}:=\underset{r\rightarrow\infty}{\lim}r^{2}G_{ab}k_{\pm}^{a}k_{\pm}^{b}$ is used in the present context. Using (37)(37)( 37 ) to set up the relation $G_{k_{\pm}k_{\pm}}=8\pi\mathcal{T}^{\pm}{}_{k_{\pm}k_{\pm}}$, where $\mathcal{T}^{\pm}{}_{k_{\pm}k_{\pm}}:=\underset{r\rightarrow\infty}{\lim}[r^{2}(T_{ab}^{\pm}\rangle_{ren}k_{\pm}^{a}k_{\pm}^{b}-\frac{1}{8\pi}\varTheta_{ab}k_{\pm}^{a}k_{\pm}^{b}]$ applies by definition, one may argue that the resulting integral relation
±$\underset{\mathcal{I}^{\pm}}{\int}N^{2}\dot{\Xi}\varepsilon_{\pm q}=\underset{\mathcal{I}^{\pm}}{\int}N^{2}[\kappa\Xi-\Xi^{2}-\sigma^{\prime}_{ab}\sigma^{\prime}{}^{ab}+\omega^{\prime}_{ab}\omega^{\prime}{}^{ab}-8\pi\mathcal{T}^{\pm}{}_{k_{\pm}k_{\pm}}]\varepsilon_{\pm q},$ (43)
constittes a semiclassical version of the null Raychaudhuri equation. This can be understood as follows:
The final term of the obtained relation in the form includes integral expressions of the form $\int\mathcal{T}^{+}{}_{k_{+}k_{+}}du$ and $\int\mathcal{T}^{-}{}_{k_{-}k_{-}}dv$, which may be viewed as subject to the averaged null energy condition (ANEC)
$\int\langle T^{+}\rangle_{ren}{}_{k_{+}k_{+}}du\geq 0,\;\int\langle T^{-}\rangle_{ren}{}_{k_{-}k_{-}}dv\geq 0.$ (44)
This condition states that along a complete null curve, any negative energy fluctuations of a quantum field must be balanced by positive energy fluctuations. Matter that violates the ANEC could, if it existed, be used to build time machines [33, 34, 64] and violate the second law of thermodynamics [78]. In contrast to most other energy conditions discussed in relativity (dominant, strong, weak, zero, etc.), however, there are no known counterexamples of the ANEC in quantum field theories (assuming that the null geodesic is achronal).
That said, given that the ANEC (44)(44)( 44 ) proves valid, it can be expected that also the conditions
±$\underset{\mathcal{I}^{\pm}}{\int}\mathcal{T}^{\pm}{}_{k_{\pm}k_{\pm}}\varepsilon_{\pm q}\geq 0$ (45)
are met. The latter can be concluded if the following two mild assumptions, which will be made below, prove to be correct: $i)$ The additional angular integration of (44)(44)( 44 ) still to be carried out has no influence on the positivity of the considered integral terms, i.e. ±$\underset{\mathcal{I}^{\pm}}{\int}\langle T^{\pm}\rangle_{ren}{}_{k_{\pm}k_{\pm}}\varepsilon_{\pm q}\geq 0$. $ii)$ The higher-curvature corrections become small enough close to past and future null infinty in the sense that ±$\underset{\mathcal{I}^{\pm}}{\int}\langle T^{\pm}\rangle_{ren}{}_{k_{\pm}k_{\pm}}\varepsilon_{\pm q}\geq\frac{1}{8\pi}\underset{\mathcal{I}^{\pm}}{\int}\varTheta{}_{k_{\pm}k_{\pm}}\varepsilon_{\pm q}$ is satisfied. By these two assumptions, it then becomes clear that the generalized ANEC (45)(45)( 45 ) is met. The validity of this relation guarantees the well-definedness of the left hand sides of (42)(42)( 42 ) and (43)(43)( 43 ) even in the event that $\mathcal{T}^{\pm}{}_{k_{\pm}k_{\pm}}<0$ applies and null rays may not be focussing. This, however, suggests that the resuting integral law can be viewed as a semiclassical generalization of the null Raychadhuri equation.
[…]
Having clarified that, one may ultimately use integral relation (43)(43)( 43 ) to set up the semiclassical quasilocal charges
$\mathfrak{Q}_{\pm}=$ $\underset{r\rightarrow\infty}{\lim}H_{h}^{Boundary}=$ (46)
$=\frac{1}{4\pi}$$\underset{\mathcal{I}^{\pm}}{\int}\left[N\mathcal{L}_{k_{\pm}}N\Xi_{\pm}+N^{2}[\varkappa_{\pm}\Xi_{\pm}-\sigma^{\prime}_{\pm ab}\sigma_{\pm}^{\prime}{}^{ab}+\omega^{\prime}_{\pm ab}\omega_{\pm}^{\prime}{}^{ab}-8\pi\mathcal{T}^{\pm}{}_{k_{\pm}k_{\pm}}]\right]\varepsilon_{\pm q},$
which take the form
$\mathfrak{Q}_{\pm}=-\frac{1}{4\pi}\underset{\mathcal{I}^{\pm}}{\int}\left[|\mathcal{\mathtt{N}}_{\pm}|^{2}+8\pi\mathcal{T}^{\pm}{}_{k_{\pm}k_{\pm}})\right]\varepsilon_{\pm q}$ (47)
in Bondi coordinates close to past and future null infinty. Alternatively, the choices $t_{\pm}^{a}=fk_{\pm}^{a}+V_{\pm}^{a}$ made in the first section lead to the related semiclassical Brown-York quasilocal charges
$\mathfrak{Q}_{f}^{\pm}=\frac{1}{4\pi}\underset{\mathcal{I}^{\pm}}{\int}\left[|\mathcal{\mathtt{N}}_{\pm}|^{2}-\frac{1}{4}D_{a}D_{b}\mathtt{N}_{\pm}^{ab}+8\pi\mathcal{T}^{\pm}{}_{k_{\pm}k_{\pm}}\right]f\varepsilon_{\pm q},$ (48)
where again some important details of the derivation are given in Appendix $A$ of this work.
Given these charges, as further clarified in more detail in the next section, the anipodal matching conditions (17)(17)( 17 ) can be readily generalized using the fact that $\mathfrak{Q}_{Y}^{\pm}\equiv Q_{Y}^{\pm}.$ This, however, prompts the question of how these generalized matching conditions can be fulfilled in the given physical setting. That question arises in particular because $\mathcal{H}$ is not modeled as a quasi-stationary event horizon in the given setting, as is usually the case, but as a dynamical horizon that changes its causal structure in the course of its evolution. The latter has an impact on the asymptotic completeness of the theory and thus on the question of whether information can be lost in the course of the evaporation process of a black hole. This will now be explained in more detail in the following, final subsection of this paper.
2.3 Dynamical Horizons and Black Hole Evaporation: New Light on Gravitational Scattering and Information Loss
The event horizon of a black hole is the future boundary of the spacetime region from which causal signals can reach $\mathcal{I}^{+}$. The mere existence of an event horizon thus presupposes knowledge of the complete Cauchy evolution of spacetime, from the very beginning to the very end. But this implies that if such a fixed, unchanging hypersurface were to form in the course of the formation and evaporation process of a black hole, it would have to exist at all times, even if the black hole has long since evaporated. The latter, however, is clearly impossible, since a black hole horizon should obviously exist no longer than the black hole singularity that created it. Once the singularity has completely evaporated, at least provided that there is no eternal black hole remnant at the end of the evaporation process, but the black hole actually evaporates in its entirety (as Hawking suggested in his early works on the subject [42, 43]), an event horizon cannot actually be part of a realistic physical model for the formation and evaporation of a black hole. Its extremely global, teleological geometric character simply does not fit the requirements of such a model.
So what if it were a dynamical horizon and not an event horizon that forms during a black hole formation and evaporation process? Would there be any substantial differences, if a quasilocal description of the process were given?
As the arguments set out in [1, 49] clearly show, there would indeed be such substantial differences. Since the inner boundary of spacetime $(\mathcal{M},g)$ was chosen to be a dynamical horizon rather than an event horizon, however, the geometric framework introduced in the first section of this work permits these arguments of the authors to be followed closely and, at the same time, to provide a quasilocal description of the process when $(\mathcal{M},g)$ is assumed to be a nonstationary black hole spacetime. Given that such a nonstationary black hole spacetime is actually being considered, the following essential differences arise if the event horizon of a black hole is replaced by a dynamical horizon:
First, in contrast to the event horizon, the dynamical horizon can be localized quasilocally and its properties are directly related to the physical processes at its location. Second, while event horizons are null hypersurfaces, dynamical horizons can either be spacelike or timelike or become null; but the latter only if they are isolated in the sense that no energy flows across them. Third, an event horizon is only one-way traversable, whereas a dynamical horizon is two-way traversable. That is, while nothing can escape from the trapped region enclosed by an event horizon to the ’outside’, causal signals can be sent across a dynamical horizon from ’inside’ to ’outside’. Thus, there is no causal obstruction for the modes that have entered the trapped region bounded by a dynamical horizon to exit again across its timelike portion.
Against this background, the following evaporation scenario can be envisioned, which is largely based on the assessment of the situation given in [1]: As the evaporation process sets in, the mass of the black hole starts to decrease and the area of the marginally trapped surfaces of the black hole goes down, thus violating the law that, classically, the area cannot decrease. This violation must apparently be caused by some flux of negative energy across the dynamical horizon into the ’interior black hole region’, which balances the positive energy flux radiated towards infinity. Heuristically, the mentioned negative energy flux should, as one would expect, be caused by quantum fluctuations leading to spontaneous (stimulated) emission, where modes are created in pairs, one of which passes the timelike portion of the dynamical horizon, while the other goes out to $\mathcal{I}^{+}$. Since the horizon of the black hole changes dynamically over time, this microscopic pair creation process has the effect that no lightlike Killing horizon is formed in the course of the evolution of the black hole, which could eventually be identified as its event horizon.
This scenario, however, can now be compared with the toy model of a Vaidya null fluid spacetime with monotonically decreasing (nonnegative) mass discussed in Appendix B of this paper. In doing so, obviously significant differences arise:
First, according to the considered toy model, there is the period $v<0$ in which the geometry of the black hole corresponds to that of an eternal Schwarschild black hole. This local spacetime geometry [51], when considered on its own right, has a null Killing horizon at $r=2M_{0}$, which would certainly be identified as the black hole’s event horizon if one were to forget the rest of the black hole’s evolution. In the time period $v>0$, this null horizon however transitions into a dynamical horizon. The mass of the black hole continues to decrease until it eventually approaches zero and the singularity at $r=0$ thus disappears. As a result, the dynamical horizon at $r(v)=2M(v)$ eventually disappears as well. Spacetime therefore becomes flat at late times. In a semiclassical context, however, it cannot be expected at all that a ’local event horizon’, i.e. a local null Killing horizon, is formed during black hole evolution.
Second, there is yet another aspect in which a fully-fledged semiclassical model should differ from the Vaidya toy model considered, even though the latter has essentially the same mass dependence at future null infinity [76]: In the Vaidya solution one has an everywhere positive energy flux that becomes locally large near the horizon, while in a semiclassical model the energy flux should be locally negative near the horizon and locally small everywhere ’outside’ the black hole (until the black hole reaches Planck dimensions).
So, there are obvious differences between the toy model used and a semiclassical model that accounts for the complete formation and evaporation of black holes. Still, there are some effects revealed by this toy model that can be expected to occur in the semiclassical regime of quantum gravity as well.
In particular, since spacetime should become flat again at late times, all field modes, including the horizon modes, can be expected to converge to the Minkowskian field modes, analogous to the Vaidya model, if the mass of the black hole actually reaches a value of zero, i.e.
$f_{m}^{+}\rightarrow f_{m}^{(0)+},\text{$g_{m}\rightarrow f_{m}^{(0)+}$}$ (49)
as $M\rightarrow 0$. As shown in Appendix B of this work, the Vaidya field modes show exactly this kind of behavior, and
$\kappa^{-1}\rightarrow 0$ (50)
is also found to apply as $M\rightarrow 0$. The latter has the effect that, by exploiting the fact that the operators
$\tilde{a}_{k}^{+}\equiv\frac{1}{\sqrt{2}}(a_{k}^{+}+b_{k}),\;\tilde{a}_{k}^{+\dagger}\equiv\frac{1}{\sqrt{2}}(a_{k}^{+\dagger}+b_{k}^{\dagger}),$ (51)
form again a CCR algebra, the decomposition (30)(30)( 30 ) of $\varphi(x)$ with respect to $\mathcal{I}^{+}\cup\mathcal{H}$ reduces to the form
$\varphi=\overset{\infty}{\underset{k=1}{\sum}}[f_{k}^{(0)+}\tilde{a}_{k}^{+\dagger}+\bar{f}_{k}^{(0)+}\tilde{a}_{k}^{+}].$ (52)
Based on the fact that the horizon modes and the other field modes take this particular form at late times, this, however, implies that a decomposition of $\varphi(x)$ can be performed with respect to purely outgoing field modes. The reason for this is that the surface gravity $\kappa$ of the black hole is not constant, but changes with time. This has the consequence that (50)(50)( 50 ) should be fulfilled as $M\rightarrow 0$, so that a purification process sets in, which ensures that eventually no thermal Hawking radiation can be detected at future null infinity.
In the semiclassical case, as already clarified, the situation is completely different. There, a decomposition of $\varphi(x)$ of the form (30)(30)( 30 ) can of course be performed, but it does not have the same meaning as for the proposed Vaidya toy model discussed in Appendix B of this paper. The reason for this is again that the horizon of spacetime is not a one-way traversable event horizon, but a two-way traversable dynamical horizon, through which matter and radiation can escape to null infinity. So there is no extended time period as in the Vaidya model where spacetime is quasi-static as the black hole forms and continuously evaporates with a dynamical horizon that is never truly isolated, i.e. never a truly non-expanding null hypersurface.
Figure 2: Graphic $(a)$ to the left shows the commonly used Penrose diagram to depict black hole evaporation and the geometric structure of the event horizon throughout the evaporation process, accounting for semiclassical quantum backreaction effects. Graphic $(b)$ on the right, on the other hand, shows a possible quantum extension of spacetime as proposed by Ashtekar in [1]. Here, the classical singularity is replaced by a transition hypersurface $\tau$, in the past of which lies a trapped region bounded in the past by a dynamical trapping horizon $T-DH$, and in the future of which lies an untrapped region bounded by a dynamical anti-trapping horizon ($AT-DH$). The graphic further displays that a purification process occurs in the future of the transition region $\tau$.
Equipped with such a horizon, the black hole evolves significantly differently than originally pictured by Hawking in [43, 44], as incoming radiative modes with negative energy can now continually affect the black hole singularity as the latter evaporates, so that there is no reason to expect a thermal Planck spectrum of emitted Hawking radiation at null infinity. Rather, the occurrence of such a spectrum can only be expected for a stationary black hole if the backreaction of the radiation on the singularity of the black hole is ignored.
Yet, since in the considered case matter and radiation can escape through the timelike portion of the black hole horizon, thereby carrying information, it can be reasonably expected that quasilocal Brown-York charges of the form (48)(48)( 48 ) can likewise reach future null infinity. These charges, as has been shown, take the form of generalized quantum analogues of the supertranslation hairs (11)(11)( 11 ) and superrotation charges (16)(16)( 16 ) first derived in [32, 46], which is why these charges should clearly be expected to satisfy the generalized antipodal matching conditions
$\mathfrak{Q}_{f,Y}^{+}=\mathfrak{Q}_{f,Y}^{-}.$ (53)
The crucial observation now is that the conditions (4952)(49-52)( 49 - 52 ) are principally strong enough to ensure that the latter is indeed the case, and furthermore to avoid any increase in entropy and thus any associated loss of quantum information along the way. This is because the asymptotic form (52)(52)( 52 ) of the scalar field suggests that radiative modes escaping to future infinity should apparently no longer be thermal in nature at the end of the evaporation process. Consequently, there appears to be no information loss in the considered case, since $i)$ the dangerous horizon mode functions take the form of Minkowskian modes near future null infinity and $ii)$ a Bogoluibov transformation of the scalar field (29)(29)( 29 ) into one of the form (52)(52)( 52 ) should lead to a unitary $S$-matrix, since only ingoing and outgoing Minkowskian field modes are being involved. Hence, it may be expected on reasonable grounds that everything that has fallen into the black hole must eventually re-emerge once the black hole is fully evaporated, albeit somewhat mangled; an expectation that proves consistent with previous results on the subject both from loop quantum gravity and superstring theory, which point to the same ultimate outcome of a black hole formation and evaporation process [4, 5, 18, 50, 60, 61, 62, 71, 73].
Summary and Conclusion
Using standard techniques from the Brown-York approach to Einstein-Hilbert gravity, quasilocal charges were derived in the present work that coincide in the large sphere limit, as was demonstrated, with conserved the supertranslation hair and superrotation charges introduced by Hawking, Perry and Strominger in [46]. Based on the results obtained, it was shown that the derived charges belong to a more general class of quasilocal Brown-York charges, which was recently derived in [54] to elucidate quasilocal corrections to the Bondi mass loss formula. By incorporating the semiclassical Einstein equations and resulting quantum backreaction effects, with a focus placed upon effects generated by massless scalar quantum fields, a generalization of the deduced soft black hole hair charges was given. Using these generalized BMS charges, which are sensitive to quantum backreactions of the spacetime geometry, a general scenario was sketched in which a black hole completely evaporates after its initial collapse due to particle creation effects, whereby a genuine event horizon never forms, but only a slowly evolving dynamical (resp. future trapping) horizon. Based on the geometric model of a Vaidya-type sandwich spacetime presented in Appendix B, it was then made clear that, according to the scenario described, it should be possible for a flux of quasilocal mass and/or radiant energy to escape through the dynamical horizon of an evaporating black hole to future null infinity. The quasilocal charges (arguably) responsible for this energy transfer were identified in this context as the quantum BMS charges derived, which generalize those of Hawking, Perry and Strominger. Ultimately, by imposing conditions on the mode functions of the scalar field, again derived with respect to the Vaidya toy model used, it was made clear that particle creation effects in nonstationary black hole spacetimes with locally defined dynamical horizons arguably cannot cause information loss at the quantum level.
That said, it may be noted that the the only evaporation scenario covered by this article is, as already emphasized, one in which the black hole fully evaporates due to the emission of Hawking radiation; thereby abandoning the possibility of the formation of a naked singularity [38, 82] or a black hole remnant in the Planck scale regime (see here, for example, [26] and references therein for information on the subject). Yet, it may also be noted that the results obtained, while showing that the singularity is naked in phases in which the black hole horizon is dynamical and thus two-way traversable, do not rule out the possibility of the formation of a black hole remnant. The main motivation for nonetheless discarding this possibility and considering only the full evaporation scenario treated above is rooted in the main objective of the present work, which is to show that the results of [42, 43, 44, 47] and [45, 46] prove to be consistent and that, therefore, Hawking’s early ideas on the subject can be reconciled with his later ideas.
The dynamical black hole spacetime considered for this purpose was modeled at the semiclassical level, which is why no singularity resolution was taken into account. This is in sharp contrast to other contemporary quantum gravity approaches to the information paradox, such as loop quantum gravity and superstring theory, where there are fundamental length scales that rule out the existence of singularities in black holes from the outset. In such more general theories of gravity, which surpass the semiclassical regime, singularity resolution would certainly have to be taken into account. Still, it can be expected on reasonable grounds that some type of similar evaporation scenario could occur as well in the theories mentioned (albeit presumably in a modified form). However, only future work in these fields will show whether the consideration of purification scenarios, as outlined in the present work, and related arguments against a loss of quantum information in evaporating black hole spacetimes will or will not actually prove useful in genuine theories of quantum gravity.
As for now, it looks like the adopted geometric framework is potentially reconcilable with the concept of holographic screens and other aspects of the theory of holography [19]. After all, both approaches are based on a similar null geometric foundation in the classical regime (though the concept of a dynamical horizon considered in this work generally differs from that of a holographic screen), and there are no strong reasons to believe that they should not prove compatible in the semiclassical regime as well. Admittedly, however, this point merits a closer examination of the subject in the future.
Ultimately, the question arises as to the experimental verification of the ideas presented. More specifically, since there are a number of approaches that suggest (or at least strive for) verification of Hawking’s results in black hole physics, e.g. through gravitational wave data or imaging of black holes or even through reallife laboratory experiments [11, 13, 66, 69, 70, 79], the question naturally arises whether purification processes and quantum information preserving effects, as presented in this paper, could also be observed or measured in the laboratory. The expectation here might be that methods similar to those likely to be used to confirm the existence of Hawking radiation could also be used to confirm evidence for the radiation purification mechanism presented in this paper. But this is, of course, pure speculation. Still, such confirmation would clearly be desirable, not least because it is generally expected that the results obtained to date in black hole thermodynamics are fundamental in the sense that they should also remain valid within the deep Planckian regime of quantum gravity. Seen in this light, any experimental evidence, however small, could ultimately shed light not only on a long-standing paradox that has attracted the greatest interest of researchers ever since its formulation almost half a century ago, but also on a theory of quantum gravity that has yet to be explored in full detail.
Appendix A
The purpose of this part of the manuscript is to further clarify how the Brown-York charges constructed in the main part of the work lead to the soft supertranslation hairs and superrotation charges constructed in [32, 46]. Considering to this end the geometric framework laid out in these works, the construction is to be carried out in retarded coordinates $(u,r,x^{A})$ near future null infinity $\mathcal{I}^{+}$ and advanced coordinates $(v,r,x^{A})$ near future past infinity $\mathcal{I}^{-}$ in the Bondi gauge. In this gauge, $g_{rr}=g_{rA}=0$ and $g_{AB}=r^{2}q_{AB}$ with $\partial_{u}q=\partial_{r}q=0$ resp. $\partial_{v}q=\partial_{r}q=0$ applies, where $q$ denotes the determinant of the induced two-metric $q_{AB}$. As a consequence, the spacetime metric takes the form
$ds^{2}=-\frac{U}{r}e^{-2\beta_{+}}du^{2}-2e^{2\beta_{+}}dudr+r^{2}q_{AB}^{+}(dx^{A}-U^{A}du)(dx^{A}-U^{A}du)$ (54)
in retarded coordinates, whereas in advanced coordinates it reads
$ds^{2}=-\frac{V}{r}e^{-2\beta_{-}}dv^{2}+2e^{2\beta_{-}}dvdr+r^{2}q_{AB}^{-}(dx^{A}-V^{A}dv)(dx^{A}-V^{A}dv).$ (55)
This form of the metric (resp. that of the associated inverse metric) then allows one to define the null vector fields $l^{a}=-\partial_{r}^{a}$, $k_{+}^{a}=e^{-2\beta}(\partial_{u}^{a}+\frac{U}{2r}\partial_{r}^{a}+U^{A}\partial_{A}^{a})$ and $k_{-}^{a}=e^{-2\beta}(\partial_{v}^{a}+\frac{V}{2r}\partial_{r}^{a}+V^{A}\partial_{A}^{a})$. Using additionally the expansions $\frac{U}{r}e^{-2\beta_{+}}=\frac{V}{r}e^{-2\beta_{-}}=1-\frac{2m}{r}+\mathscr{O}(r^{-2})$, $e^{2\beta_{\pm}}=1-\frac{C_{AB}^{\pm}C_{\pm}^{AB}}{32r^{2}}+\mathscr{O}(r^{-3})$, $U^{A}=u_{A}+\mathscr{O}(r^{-4})$ with $u_{A}:=-\frac{1}{2}D^{B}C_{+\>B}^{A}-\frac{1}{r^{3}}[\frac{2}{3}\mathcal{\mathtt{N}}_{A}+\frac{2u}{3}D_{A}m-\frac{1}{16}D_{A}(C_{CD}^{+}C_{+}^{CD})]$, $v_{A}:=-\frac{1}{2}D^{B}C_{-\>B}^{A}-\frac{1}{r^{3}}[\frac{2}{3}\mathcal{\mathtt{N}}_{A}+\frac{2v}{3}D_{A}m-\frac{1}{16}D_{A}(C_{CD}^{-}C_{-}^{CD})]$ and $q_{AB}^{\pm}=(1+\frac{1}{4r^{2}}C_{CD}^{\pm}C_{\pm}^{CD})\gamma_{AB}+\frac{1}{r}C_{AB}^{\pm}+\mathscr{O}(r^{-3})$, where $\mathcal{\mathtt{N}}_{A}$ is the angular momentum aspect and $\gamma_{AB}$ is a the induced metric on the sphere, the metric reduces to the form
$ds^{2}=-(1-\frac{2m}{r})du^{2}-2dudr+u_{A}dudx^{A}+r^{2}q_{AB}^{+}dx^{A}dx^{B}$ (56)
resp.
$ds^{2}=-(1-\frac{2m}{r})dv^{2}+2dvdr+v_{A}dvdx^{A}+r^{2}q_{AB}^{-}dx^{A}dx^{B}$ (57)
in the introduced coordinates. To identify the lapse function $N$ of the geometry, it proves convenient to introduce new coordinates $u=t-r$ and $v=t+r$. In these coordinates (51)(51)( 51 ) and (52)(52)( 52 ) can be rewritten in the form
$ds^{2}$ $=-(1-\frac{2m}{r})dt^{2}-\frac{4m}{r}dtdr+(1+\frac{2m}{r})dr^{2}+u_{A}(dt-dr)dx^{A}$ (58)
$+r^{2}q_{AB}^{+}dx^{A}dx^{B}$
resp.
$ds^{2}$ $=-(1-\frac{2m}{r})dt^{2}+\frac{4m}{r}dtdr+(1+\frac{2m}{r})dr^{2}+v_{A}(dt+dr)dx^{A}$ (59)
$+r^{2}q_{AB}^{-}dx^{A}dx^{B},$
thereby allowing one to identify $N(r)=\sqrt{1-\frac{2m}{r}}$ as the lapse function to the geometry. That said, a possible choice for the null vector fields of the geometry in the given coordinate setting is $l^{a}=-\partial_{r}^{a}$, $k_{+}^{a}=(\partial_{u}^{a}-\frac{N}{2}\partial_{r}^{a})$ and $k_{-}^{a}=(\partial_{v}^{a}+\frac{N}{2}\partial_{r}^{a})$. The latter, however, implies that $\partial_{u}^{a}=k_{+}^{a}-\frac{N}{2}l^{a}$ and $\partial_{v}^{a}=k_{+}^{a}+\frac{N}{2}l^{a}$.
Given all that, the stress-energy tensor of spacetime shall be assumed to satisfy the following fall-off conditions as $r\rightarrow\infty$: $T_{uu}(u,r,x^{A})=\frac{\mathfrak{T}{}_{uu}(u,x^{A})}{r^{2}}+\mathscr{O}(r^{-3})$, $T_{uA}(u,r,x^{A})=\frac{\mathfrak{T}_{uA}(u,x^{A})}{r^{2}}+\mathscr{O}(r^{-3})$, $T_{AB}(u,r,x^{A})=\frac{\mathfrak{T}(u,x^{A})\gamma_{AB}}{r}+\mathscr{O}(r^{-2})$, $T_{ur}(u,r,x^{A})=\mathscr{O}(r^{-4})$, $T_{rr}(u,r,x^{A})=\mathscr{O}(r^{-4})$, $T_{rA}(u,r,x^{A})=\mathscr{O}(r^{-3})$ resp. $T_{vv}(v,r,x^{A})=\frac{\mathfrak{T}_{vv}(v,x^{A})}{r^{2}}+\mathscr{O}(r^{-3})$, $T_{vA}(v,r,x^{A})=\frac{\mathfrak{T}_{vA}(v,x^{A})}{r^{2}}+\mathscr{O}(r^{-3})$, $T_{AB}(v,r,x^{A})=\frac{\mathfrak{T}(v,x^{A})\gamma_{AB}}{r}+\mathscr{O}(r^{-2})$, $T_{vr}(v,r,x^{A})=\mathscr{O}(r^{-4})$, $T_{rr}(v,r,x^{A})=\mathscr{O}(r^{-4})$, $T_{rA}(v,r,x^{A})=\mathscr{O}(r^{-3})$. These conditions yield
$T_{kk}^{+}=$ $\underset{r\rightarrow\infty}{\lim}r^{2}T_{ab}k_{+}^{a}k_{+}^{b}=\underset{r\rightarrow\infty}{\lim}r^{2}T_{uu}=\mathfrak{T}_{uu},$ (60)
$T_{kk}^{-}=$ $\underset{r\rightarrow\infty}{\lim}r^{2}T_{ab}k_{-}^{a}k_{-}^{b}=\underset{r\rightarrow\infty}{\lim}r^{2}T_{vv}=\mathfrak{T}_{vv}.$
Moreover, introducing next the past and future news tensors $\mathcal{\mathtt{N}}_{AB}^{+}=\partial_{u}C_{AB}^{+}$ and $\mathcal{\mathtt{N}}_{AB}^{-}=\partial_{v}C_{AB}^{-}$, where $C_{AB}^{\pm}=-2D_{A}D_{B}C|_{\mathcal{I}_{\mp}^{\pm}}+\gamma_{AB}D^{2}C|_{\mathcal{I}_{\mp}^{\pm}}$ with $C^{+}=C^{+}(u,x^{A})$ and $C^{-}=C^{-}(v,x^{A})$ shall apply by definition, let it be be required - along the same lines as in [46] - that near the past and future boundaries of $\mathcal{I}^{\pm}$ and $\mathcal{I}_{\mp}^{\pm}$, the news falls off faster than $\frac{1}{|u|}$ resp. $\frac{1}{|v|}$ and that the angular momentum aspect $\mathcal{\mathtt{N}}_{A}$ approaches a finite one-form on $\mathbb{S}_{2}$. This then makes it clear that the bulk is not required to be nearly flat far from the boundary, but can have any geometric structure whatsoever.
Using the above, it then becomes clear that the quasilocal charges depicted in (10)(10)( 10 ) take the form
$Q_{+}$ $=\frac{1}{4\pi}\underset{\mathcal{I}^{+}}{\int}\left[\frac{1}{8}\mathcal{\mathtt{N}}_{AB}^{+}\mathcal{\mathtt{N}}_{+}^{AB}+8\pi\mathfrak{T}_{uu}\right]\varepsilon_{+\gamma},$ (61)
$Q_{-}$ $=\frac{1}{4\pi}\underset{\mathcal{I}^{-}}{\int}\left[\frac{1}{8}\mathcal{\mathtt{N}}_{AB}^{-}\mathcal{\mathtt{N}}_{-}^{AB}+8\pi\mathfrak{T}_{vv}\right]\varepsilon_{-\gamma}.$
thus leading one to conclude that the expressions obtained coincide - up to a total divergence term equal to zero - with the families of charges defined in [^1] To be precise, the charges occur in the special case $f=1$. [46]. Moreover, taking relations (1416)(14-16)( 14 - 16 ) into account, one obtains the further results
$J_{\mathcal{I}_{-}^{+}}^{\phi}=\frac{1}{4\pi}\underset{\mathcal{I}_{-}^{+}}{\int}\Omega_{A}\phi^{A}\omega_{\gamma}$ $=\frac{1}{8\pi}\underset{\mathcal{I}_{-}^{+}}{\int}\mathcal{\mathtt{N}}_{A}Y^{A}\omega_{\gamma}=Q_{Y}^{+},$ (62)
$J_{\mathcal{I}_{+}^{-}}^{\phi}=\frac{1}{4\pi}\underset{\mathcal{I}_{+}^{-}}{\int}\Omega_{A}\phi^{A}\omega_{\gamma}$ $=\frac{1}{8\pi}\underset{\mathcal{I}_{+}^{-}}{\int}\mathcal{\mathtt{N}}_{A}Y^{A}\omega_{\gamma}=Q_{Y}^{-},$
provided that $\mathcal{\mathtt{N}}^{A}$ denotes the angular momentum aspect and $Y^{A}$ is a conformal Killing vector field on the two-sphere given by $Y^{A}=D^{A}\chi+\varepsilon^{AB}D_{B}\kappa$ with $\chi$ and $\kappa$ representing solutions of the differential relation $(D^{2}+2)\chi=(D^{2}+2)\kappa=0$ for $D^{2}:=D_{A}D^{A}.$
Considering now as in [46] a supertranslation generating diffeomorphism vector fields of the form $\zeta_{+}^{a}=f\partial_{u}^{a}-\frac{1}{r}D^{A}f\partial_{A}^{a}+\frac{1}{2}D^{2}f\partial_{r}^{a}$ with $\partial_{u}f=\partial_{r}f=0$ and $\zeta_{-}^{a}=f\partial_{v}^{a}-\frac{1}{r}D^{A}f\partial_{A}^{a}+\frac{1}{2}D^{2}f\partial_{r}^{a}$ with $\partial_{v}f=\partial_{r}f=0$, the same line of reasoning that led to the form of (61)(61)( 61 ) can be used to conclude that Brown-York charges (11)(11)( 11 ) reduce to the form
$Q_{f}^{+}$ $=\frac{1}{4\pi}\underset{\mathcal{I}^{+}}{\int}\left[\frac{1}{8}\mathtt{N}_{AB}^{+}\mathtt{N}_{+}^{AB}-\frac{1}{4}D_{A}D_{B}\mathtt{N}_{+}^{AB}+8\pi\mathfrak{T}_{uu}\right]f\varepsilon_{+\gamma},$ (63)
$Q_{f}^{-}$ $=\frac{1}{4\pi}\underset{\mathcal{I}^{-}}{\int}\left[\frac{1}{8}\mathtt{N}_{AB}^{-}\mathtt{N}_{-}^{AB}-\frac{1}{4}D_{A}D_{B}\mathtt{N}_{-}^{AB}+8\pi\mathfrak{T}_{vv}\right]f\varepsilon_{-\gamma},$
if $t^{a}\equiv\zeta_{\pm}^{a}$ is chosen as the time evolution vector field of spacetime and the boundary of the latter is shifted to infinity in the large sphere limit, where it has been used that $|\mathcal{\mathtt{N}}_{\pm}|^{2}=\frac{1}{8}\mathtt{N}_{AB}^{\pm}\mathtt{N}_{\pm}^{AB}$. This shows that the conserved supertranslation hair and superrotation charge expressions arise naturally in the given coordinate setting from the quasilocal Brown-York charges defined in section one and two of this manuscript.
That said, the antipodal matching conditions (17)(17)( 17 ) are found to be equivalent to the Lorentz and CPT invariant matching conditions
$C|_{\mathcal{I}_{-}^{+}}=C|_{\mathcal{I}_{+}^{-}},\>m|_{\mathcal{I}_{-}^{+}}=m|_{\mathcal{I}_{+}^{-}},\>N^{A}|_{\mathcal{I}_{-}^{+}}=N^{A}|_{\mathcal{I}_{+}^{-}}$ (64)
formulated in [46], thereby proving consistency with the Brown-York approach adopted in this work.
Eventually, given the transition $g_{ab}\rightarrow g_{ab}+L_{\zeta}g_{ab}$, where the variations $L_{\zeta}g_{ab}$ around the background metric $g_{ab}$ are typically assumed to obey the linearized Einstein equations, it was shown in [46] that one can now use the foregoing to $i)$ ’implant’ supertranslation hairs in black hole spacetimes and $ii)$ construct horizon charges whose explicit form depends on the choice of coordinates used to extend the supertranslations in from the boundary to the horizon.
Such charges are constructed explicitly in [46]. It is worth noting that the above results show that the construction of such horizon charges or supertranslation hair in black hole spacetimes is, in principle, compatible with the Brown-York framework used in the present work, even though said horizon charges have not been considered in the main body of the manuscript.
Instead quasilocal charge resulting from the semiclassical Einstein equations (37)(37)( 37 ) were calculated, where the latter may be rewritten in the form
$E_{\;b}^{a}=G_{\;b}^{a}-8\pi\mathcal{T}_{\;b}^{a}=0.$ (65)
To perform a semiclassical extension of the standard Bondi-Sachs formalism (see here, for instance, [59] and references therein for an introduction to the subject), it shall be assumed in the following that a renormalization scheme is used according to which the renormalized expectation value of the stress-energy tensor (in combination with all resulting higher-order curvature corrections) is covariantly conserved. Under this assumption, the contracted Bianchi identity
$\nabla_{a}E_{\;b}^{a}=0$ (66)
is satisfied in the same way as in the classic Bondi-Sachs approach. In consequence, the hypersurface and evolution equations
$E_{\;a}^{u}=0,\;E_{AB}-\frac{1}{2}q_{AB}E=0,$ (67)
resp.
$E_{\;a}^{v}=0,\;E_{AB}-\frac{1}{2}q_{AB}E=0,$ (68)
can be designated close to future and past null infinity, where $E=q^{CD}E_{CD}$ with $C,D=2,3$ shall apply by definition. For the purpose of integrating these main equations, in principle, the same steps as in the standard classical Bondi-Sachs formalism can be taken. In particular, semiclassical metric data similar to those in (54)(54)( 54 ) and (55)(55)( 55 ) can be prescribed as a basis for the integration of (67)(67)( 67 ) and (68)(68)( 68 ), which in turn leads to line elements of the form (56)(56)( 56 ) and (57)(57)( 57 ) if the same fall-off conditions as at the beginning of this Appendix are imposed. Eventually, if the same fall-off conditions are invoked for the renormalized expectation value of the stress-energy tensor operator as in the above case for the classical energy-momentum tensor, i.e., if $T_{ab}(x)$ and $\mathcal{T}_{ab}(x)$ exhibit the same asymptotic fall-off behavior, the formalism used yields semiclassical analogues of the quasilocal Brown-York charges (10)(10)( 10 ) and (11)(11)( 11 ), the explicit form of which is depicted in (47)(47)( 47 ) and (48)(48)( 48 ).
Appendix B
As it proves to be too difficult to model a black hole formation and evaporation process with semiclassical methods, an obvious alternative is to consider a manageable spherically symmetric model that covers at least some essential properties of the geometry of an evaporating black hole. Such a model is given by the Vaidya geometry
$ds^{2}=-(1-\frac{2M}{r})dv^{2}+2dvdr+r^{2}d\Omega^{2},$ (69)
which describes the gravitational field of a null fluid with time-dependent mass $M(v)$ and a stress-energy tensor of the form
$T_{\;b}^{a}=\frac{\dot{M}}{4\pi r}k^{a}k_{b}.$ (70)
To describe the evaporation part of the evolution of the black hole, it makes sense to consider the mass function
[…]
with $v_{0}=\frac{3c}{M_{0}^{3}}$. Accordingly, the evaporation process take places in the time interval $[0,v_{0}]$, whereby, depending on the concrete value of the constant $c$, it may be assumed that the geometry of the black hole evolves slowly. The particular choice made for the mass function in (71)(71)( 71 ) give rises to the evolution equation
$\frac{dM}{dv}=-\frac{c}{M^{2}}$ (72)
in the time period $[0,v_{0}]$, in full agreement with Stefan-Boltzmann’s law. The focus is thus solely placed on the evaporation process, whereas the formation process of the black hole is completely disregarded. The black hole has simply always existed in the period $v<0$ with a distributional stress-energy tensor of the form [8, 9]
$T_{\;b}^{a}=-M_{0}\delta^{(3)}(x)v^{a}v_{b},$ (73)
where $v^{a}$ is a timelike vector field ’outside’ the $r=2M_{0}$-region.
As can be seen, the geometry has a Kodama vector field $\xi^{a}=\partial_{v}^{a}$ that constitutes a genuine Killing vector field at early and late times [57], where the latter can be identified as a hidden Killing field in the sense of [52]. As may further be noted, the dynamical horizon of the geometry lies at $r_{\mathcal{H}}(v)=2M(v)$. Since the black hole evaporates very slowly, it can be concluded that $M(v)\approx M_{0}$ at the beginning of the evaporation period, thereby implying that the dynamical horizon $\mathcal{H}$ behaves in that time period almost like a quasi-static black hole horizon with associated constant radial parameter $r_{0}=2M_{0}$, which can actually be identified as an isolated horizon in the sense of Ashtekar et al. [2, 3, 7]. Still, this horizon is not a genuine event horizon. It is merely a local event horizon for the local black hole spacetime, which exists in the time span $v\in(-\infty,0]$; for further details on local spacetimes, see [51].
In any case, given the above geometric setting, let the wave equation (28)(28)( 28 ) be considered next using the identity $\boxempty=\frac{1}{\sqrt{-g}}\partial_{a}(\sqrt{-g}g^{ab}\partial_{b})$. After separating the variables here as
$\varphi_{\omega lm}(v,r,\theta,\phi)=\frac{1}{r}\rho(v,r;\omega)Y_{lm}(\theta,\phi),$ (74)
one finds that the equation splits up in two parts: a standard angular part, which can straightforwardly be solved and thus will not be of much interest in the following, and a radial part
$(1-\frac{2M}{r})\frac{\partial^{2}\rho}{\partial r^{2}}+2\frac{\partial^{2}\rho}{\partial r\partial v}+\frac{2M}{r^{2}}\frac{\partial\rho}{\partial r}-[\frac{2M}{r^{3}}+\frac{l(l+1)}{r^{2}}]\rho=0,$ (75)
which will be the main focus of attention in what follows. Following closely the arguments used in [56, 81], this radial part can be rewritten by introducing the tortoise coordinate $r_{*}=r+\frac{1}{2\kappa}\ln|\frac{r-r_{\mathcal{H}}}{r_{\mathcal{H}}}|$ in the form
$A\frac{\partial^{2}\rho}{\partial r_{*}^{2}}+2\frac{\partial^{2}\rho}{\partial r_{*}\partial v}+B\frac{\partial\rho}{\partial r_{*}}-C[\frac{2M}{r^{3}}+\frac{l(l+1)}{r^{2}}]\rho=0,$ (76)
provided that the definitions $A(v,r)=\frac{(2\kappa(r-r_{\mathcal{H}})+1)(r-2M)-2r\dot{r}_{\mathcal{H}}}{2\kappa(r-r_{\mathcal{H}})}$, $B(v,r)=\frac{2r\dot{r}_{\mathcal{H}}-r+2M}{r(r-r_{\mathcal{H}})(2\kappa(r-r_{\mathcal{H}})+1)}+\frac{2M}{r^{2}}$ and $C(v,r)=\frac{2\kappa(r-r_{\mathcal{H}})}{2\kappa(r-r_{\mathcal{H}})+1}$ are used in the present context. Supposing that $\underset{r\rightarrow r_{\mathcal{H}}}{\lim}A(v,r)\rightarrow 1$, one then finds
$\underset{r\rightarrow r_{\mathcal{H}}}{\lim}(r-2M+2r\dot{r}_{\mathcal{H}})\rightarrow 0,$ (77)
whence
$r_{\mathcal{H}}=\frac{2M}{1-2\dot{r}_{\mathcal{H}}},\;\kappa=\frac{1-2\dot{r}_{\mathcal{H}}}{4M}.$ (78)
For small $\dot{r}_{\mathcal{H}}$ (and hence $\dot{M}$), it can thus be concluded that
$r_{\mathcal{H}}\approx 2M(1+4\dot{M}),\;\kappa\approx\frac{1-4\dot{M}}{4M}$ (79)
in agreement with the results of Babinot [10] and York [80]. At the very beginning of the evaporation process, one may safely assume that $\dot{M}\approx 0$, which, however, allows one to conclude that
$r_{\mathcal{H}}\approx 2M,\;\kappa_{0}\approx\frac{1}{4M}$ (80)
applies in this particular period of time. As a result, equation (76)(76)( 76 ) reduces to the form
$\frac{\partial^{2}\rho}{\partial r_{*}^{2}}+2\frac{\partial^{2}\rho}{\partial r_{*}\partial v}+2\kappa_{0}\frac{\partial\rho}{\partial r_{*}}=0,$ (81)
as follows from the fact that $\underset{r\rightarrow r_{\mathcal{H}}}{\lim}A(r,v)\rightarrow 1$, $\underset{r\rightarrow r_{\mathcal{H}}}{\lim}B(r,v)\rightarrow 2\kappa_{0}(v)$ and $\underset{r\rightarrow r_{\mathcal{H}}}{\lim}C(r,v)\rightarrow 0$.
The differential equation thus obtained can be solved. It has the two independent solutions
$\rho_{\omega}^{in}=e^{-i\omega v+\mathcal{K}_{0}},\;\rho_{\omega}^{out}=e^{-i\omega(v-2r_{*})+\mathcal{K}_{0}},$ (82)
[…]
As may be noted, the part $e^{-i\omega(v-2r_{*})}=e^{-i\omega(v-2r)}(r-r_{\mathcal{H}})^{\frac{i\omega}{\kappa_{0}}}$ is not analytic at the horizon $r_{\mathcal{H}}$. Following [81], however, the mentioned part can be analytically continued from $r>r_{\mathcal{H}}$ to $r\leq r_{\mathcal{H}}$ such that
$r-r_{\mathcal{H}}$ $\rightarrow|r-r_{\mathcal{H}}|e^{-i\pi}=(r-r_{\mathcal{H}})e^{-i\pi},$ (83)
$\rho_{\omega}^{out}$ $\rightarrow\tilde{\rho}_{\omega}^{out}=e^{\frac{\pi\omega}{\kappa_{0}}}\rho_{\omega}^{out}.$
Thus, the relative scattering probability of the outgoing wave at the horizon is
$|\frac{\rho_{\omega}^{out}}{\tilde{\rho}_{\omega}^{out}}|^{2}=e^{-\frac{2\pi\omega}{\kappa_{0}}},$ (84)
thereby implying that the spectrum of the Hawking radiation from the horizon is
$N_{\omega}=\frac{1}{e^{\frac{2\pi\omega_{k}}{\kappa_{0}}}-1}.$ (85)
From this, in turn, it can be concluded that the black hole emits thermal Hawking radiation of temperature
$T=\frac{\kappa_{0}}{2\pi}.$ (86)
The field modes (83)(83)( 83 ) and (84)(84)( 84 ) can therefore be used to perform a splitting of the form (30)(30)( 30 ) of $\varphi_{\omega lm}(v,r,\theta,\phi)$ (possibly after smearing them out with test functions with compact support at the dynamical horizon $\mathcal{H}$ and at null infinity). Still, this splitting cannot simply be regarded as a conventional splitting into future null infinity field modes and horizon modes. To see why, one may recall that in the local flat Minkwoskian part of the spacetime occurring for $v>v_{0}$ relations (75)(75)( 75 ) and (76)(76)( 76 ) reduce to the form
$\frac{\partial^{2}\rho}{\partial r^{2}}+2\frac{\partial^{2}\rho}{\partial r\partial v}=0,$ (87)
thereby giving rise to two independent solutions of the form
$\rho_{0\omega}^{in}=e^{-i\omega v},\;\rho_{0\omega}^{out}=e^{-i\omega(v-2r)}.$ (88)
In the Schwarzschild case occurring for $v<0$, on the other hand, differential relation (76)(76)( 76 ) becomes
$\frac{\partial^{2}\rho}{\partial r_{0*}^{2}}+2\frac{\partial^{2}\rho}{\partial r_{0*}\partial v}+2\kappa_{0}\frac{\partial\rho}{\partial r_{0*}}=0,$ (89)
for small variations of the mass parameter (i.e. for $\dot{M}(v)\approx 0$ and $M(v)\approx M_{0}\approx const.$), where $r_{0*}=r+2M_{0}\ln|\frac{r-2M_{0}}{2M_{0}}|$ applies by definition. The solutions of this equation read
$\rho_{\omega}^{in}=e^{-i\omega v+\mathcal{K}_{0}},\;\rho_{\omega}^{out}=e^{-i\omega(v-2r_{0*})+\mathcal{K}_{0}},\;\tilde{\rho}_{\omega}^{out}=e^{\frac{\pi\omega}{\kappa_{0}}}\rho_{\omega}^{out},$ (90)
and thus coincide exactly with (83)(83)( 83 ) and (84)(84)( 84 ) for $M(v)\approx M_{0}$. In this case, the first law of black hole mechanics [12] dictates that $\kappa_{0}=const.$ applies, so that it can be concluded that $\mathcal{K}_{0}(v)=\kappa_{0}v$ applies in the time span $v<0$. This changes for $v\in[0,\infty)$; a period in which it can be observed that $\kappa_{0}^{-1}\rightarrow 0,\;\mathcal{K}_{0}\rightarrow 0$ as $M(v)\rightarrow 0$. Consequently, however, it can be concluded that the dynamical mode functions (83)(83)( 83 ) and (84)(84)( 84 ) at late times take the form
$\rho_{\omega}^{in}\rightarrow\rho_{0\omega}^{in},\;\rho_{\omega}^{out}\rightarrow\rho_{0\omega}^{out},\;\tilde{\rho}_{\omega}^{out}\rightarrow\rho_{0\omega}^{out}$ (91)
and thus coincide with the Minkowskian mode functions (88)(88)( 88 ) for $v\in[v_{0},\infty)$. As a result, the radiation escaping to future null infinity appears no longer to be thermal in nature at the end of the evaporation process. It is therefore not expected, according to the proposed toy model, that a thermal spectrum of the form (35)(35)( 35 ) or (85)(85)( 85 ) occurs at late epochs of the evolution of spacetime.
That said, let it ultimately be noted that comparable results can also be derived for a supertranslated Vaidya spacetime, as considered in particular in [29], if a mass parameter of the form (71)(71)( 71 ) is considered. Despite the fact that the expression for the surface gravity takes a different form in said case (see here equation 3.123.123.12 on page 101010), still there is full confirmation of the results obtained in the present work. Thus, once again, there is good agreement with the previously known literature on the subject.
Acknowledgements:
I want to thank Felix Wilkens for his support in preparing the images depicted in Figures 111 and 222 of the paper. Moreover, I want to acknowledge the UAS Technikum Wien Library for financial support through its Open Access Funding Program.